### v2

# Interpolation in Weighted Projective Spaces

Abstract.
Over an algebraically closed field, the double point interpolation problem asks for the vector space dimension of the projective hypersurfaces of degree $d$ singular at a given set of points. After being open for 90 years, a series of papers by J. Alexander and A. Hirschowitz in 1992–1995 settled this question in what is referred to as the Alexander-Hirschowitz theorem. In this paper we primarily use commutative algebra to lay the groundwork necessary to prove analogous statements in the weighted projective space, a natural generalization of the projective space. We prove the Hilbert function of general simple points in any $n$-dimensional weighted projective space exhibits the expected behavior. We also introduce an inductive procedure for weighted projective space, similar to that originally due to A. Terracini from 1915, to demonstrate an example of a weighted projective plane where the analogue of the Alexander-Hirschowitz theorem holds without exceptions and prove our example is the only such plane. Furthermore, Terracini’s lemma regarding secant varieties is adapted to give an interpolation bound for an infinite family of weighted projective planes.
The author received support from the National Science Foundation grant DMS-210225.
Mathematics Subject Classification: 13A02, 13F20, 14N05.
Key Words: Interpolation, Hilbert Function, Weighted Projective Space, Non-standard Grading.
1. Introduction
Let $k$ be an algebraically closed field. Consider the vector space of homogeneous polynomials of degree $d$ in $n+1$ variables having singularities at $r$ general points $p_{1},\ldots,p_{r}$ of the projective space ${\mathbb{P}}^{n}_{k}$. For a polynomial $f$ to be a member of this vector space it must satisfy $r(n+1)$ linear conditions on its coefficients imposed by the vanishing of all partial derivatives of $f$ at $p_{1},\ldots,p_{r}$. It is natural to inquire whether these conditions are independent. This problem was considered by A. Hirschowitz and J. E. Alexander in a series of papers [Hir85, Ale88, AH92a, AH92b, AH95] and finally settled in the form stated below, where $H_{d}(-)$ denotes the dimension of the vector subspace of homogeneous forms of degree $d$ of the input.
Theorem 1.1 (Alexander-Hirschowitz).
Let $n,d$ be positive integers. Let $X$ be a general set of $r$ double points in ${\mathbb{P}}^{n}_{k}$ and $I={I_{X}}$ be its defining ideal. Then
$H_{d}(R/I)=\min\left\{r(n+1),{{n+d}\choose{d}}\right\}$ (1.1)
with exceptions:
  • $d=2$ and $2\leq r\leq n$;
  • $d=3$, $n=4$ and $r=7$; and
  • $d=4$, $2\leq n\leq 4$ and $r={n+2\choose 2}-1$.
We say that a set of double points $X$ has expected dimension in degree $d$ if its ideal $I$ satisfies Equation 1.1. Theorem 1.1 has two equivalent formulations in terms of the dimension of higher secant varieties to the Veronese embedding of ${\mathbb{P}}^{n}$ and the Big Waring Rank problem ([HM21, Appendix A]) as well as applications in numerical algebraic geometry [DREE+23], deep polynomial neural networks [KTB19], and algebraic statistics [Cer17, Section 6.2]. The original proof of Theorem 1.1 is based on a degeneration technique called the “la méthode d’Horace différentialle”, or méthode d’Horace for short, which is based on an earlier degeneration technique by A. Terracini (Theorem 6.2). Méthode d’Horace was later simplified by K. A. Chandler [Cha01, Cha02]. Further expositions and surveys on the Alexander-Hirschowitz theorem and related problems can be found in [BO08, HM21].
Our main goal is to give an analogous statement to Theorem 1.1 in the weighted projective space denoted ${\mathbb{P}}(a_{0},\ldots,a_{n})$, where each $a_{i}$ is a positive integer. This space, surveyed in [Dol82, BR86], arises naturally in contexts such as toric geometry [RT11], algebraic topology [BFNR13], and $K$-theory [Amr94]. When $a_{i}=1$ for all $i$, one can use Definition 2.1 to recover the usual or straight projective space. The author’s main approach has been to modify existing techniques from the usual projective space to fit this new setting. Such adaptions require careful, and often non-trivial, considerations of details. This is best exemplified in Appendix B.
In Section 2 we introduce the weighted projective space, its basic properties and the defining ideal of points. Proposition 2.8 shows that unlike the projective space, even a general point in the weighted projective plane need not be generated by a complete intersection ideal; a fact which is critical to computing $I$ in Theorem 1.1. However, having a weight of $1$ forces the ideals of general points to be complete intersections, Corollary 2.9, hence allowing for easy calculations of their ideals.
Section 3 concerns an analogue to Theorem 1.1 in the weighted projective line and shows considering this property passes a litmus test, it holds in dimension $1$. Its main result is Proposition 3.2 which we now state.
Proposition 3.2.
A general set of points with mixed multiplicity in ${\mathbb{P}}(a,b)$, where $\gcd(a,b)=1$, has expected dimension in degree $d$ for all $d\in{\mathbb{N}}$.
Here, mixed multiplicity refers to having different orders of singularity at the points. While for $n=1$ Theorem 1.1 does not require the assumption that the points are general, in contrast, our result excludes the point $[0:1]$, hence the hypothesis of Proposition 3.2 requires general points.
[…]
Theorem 4.1.
Let $X\subseteq{\mathbb{P}}(a_{0},a_{1},\ldots,a_{n})$ be a set of $r$ general simple points with defining ideal $I_{X}$. Then for all $d\in{\mathbb{N}}$ we have
$H_{d}(S/I_{X})=\min\{r,H_{d}(S)\}.$
The proof of Theorem 4.1 in the usual projective space is accessible to an advanced undergraduate student; see [Hui18]. But a proof in the setting above is non-trivial and requires tempering with parameter spaces.
In Section 5 we study sets of double points in weighted projective planes of the form ${\mathbb{P}}(1,b,c)$. The main novelty of this section is taking a different view point towards the problem: instead of studying interpolation in a fixed space, we start with a fixed size for a set of general double points, say $r$, and find examples of spaces in which $r$ general double points have expected dimension in every degree $d$. Proposition 5.4 proves a necessary condition for this to hold.
Proposition 5.4.
If $r$ general double points in ${\mathbb{P}}(1,b,c)$ have expected dimension in every degree, then $c<(r+1)b$ for $b\neq 1$ and $c\leq r+1$ when $b=1$.
The main result of Section 5 is a partial converse to Proposition 5.4.
Proposition 5.8.
Let $r\in{\mathbb{N}}$. If $rb<c<(r+1)b$ then a set of $r$ general double points in ${\mathbb{P}}(1,b,c)$ has expected dimension in every degree $d$.
The entirety of Section 6 is dedicated to proving an Alexander–Hirschowitz theorem in ${\mathbb{P}}(1,2,3)$.
Theorem 6.1.
Let $X\subseteq{\mathbb{P}}(1,2,3)$ be a set of $r$ general double points. Then $X$ has expected dimension in every degree $d$ with no exceptions. Moreover ${\mathbb{P}}(1,2,3)$ is the only weighted projective space where no exceptions occur.
The most striking part of Theorem 6.1 is having no exceptions which is not the case even in the usual ${\mathbb{P}}^{2}$. Indeed, Theorem 7.11 shows ${\mathbb{P}}(1,2,3)$ is the only such plane. The main inductive procedure of the proof, Theorem 6.2, is the predecessor to méthode d’Horace; a specialization technique due to Terracini from 1915, [Ter15a]. In Theorem 6.2 we observe Terracini’s method can be generalized to ${\mathbb{P}}(1,\ldots,a_{n})$. The difference in numerology is key; ${\mathbb{P}}_{k}^{n}$ contains only one hyperplane up to isomorphism. But in ${\mathbb{P}}(1,\ldots,a_{n})$ there are $n+1$ hyperplanes with different numerical characteristics.
Theorem 1.1 provides a finite list of exceptions in ${\mathbb{P}}^{2}$; $2$ points in degree $2$ and $5$ points in degree $4$. In Example 7.2 we show that general double points, even in a “nice” space such as ${\mathbb{P}}(1,b,c)$ can have erratic behavior. Nonetheless, Theorem 7.7 gives a linear bound, in terms of the weights, on the degree $d$, after which a set of points of any size expected dimension.
Theorem 7.7.
Let $X\subseteq{\mathbb{P}}(1,b,c)$ be set of $r$ general double points. Then $X$ has expected dimension in degree $d$ if $d\geq 10c$ or if $d\geq 6c$ and $\left\lfloor\frac{2c}{b}\right\rfloor\geq 5$.
The case of the projective plane of Theorem 1.1 had been previously elucidated by Terracini in [Ter15b] and by Platini in [Pal03]. Our proof of Theorem 7.7 mimics Terracini’s approach. However the bound in our result depends on the weights. This is a case where the numerics of weighted projective space work to our disadvantage: as the weights increase, one has more cases to check.
The essential ingredient in proving Theorem 7.7 is the equivalence of double point interpolation and computing the dimension of the higher secant variety to the Veronese embedding of projective space. This is made precise in Theorem 7.5, the proof of which is relegated to Appendix B. The main requisite in proving this equivalence is due to Lasker Proposition B.5.
We end the introduction highlighting three major themes of this paper: the divergence of notions from the usual projective space when they are considered in the weighted projective space, the difference in numerology and the importance of a weight of $1$. For example, in Definition 7.4, one observes when $a_{i}\neq 1$ for all $i$, the $d$-th Veronese map is not defined on all of ${\mathbb{P}}(a_{0},\ldots,a_{n})$, hence it is not an embedding; a fact crucial to the analysis of its secant variety. Other examples of this divergence are the defining ideal of a simple point not being complete intersection, addressed earlier, or stabilization of the Hilbert function at the expected number of independent conditions; see Remark 5.2 (2). In all cases, introducing a single weight of $1$ controls this divergence to some extent. All specialization techniques used for the interpolation problem rely heavily on numerical conditions [AH95, Cha01, BO08, Pos12]. While for a standard graded polynomial ring the Hilbert function can be computed using a binomial coefficient, no closed formula exists in the non-standard graded setting beyond $2$ variables. Therefore applying degeneration techniques becomes very tricky. The remarkable fact that the Hilbert function for the coordinate ring of ${\mathbb{P}}(1,2,3)$ happens to have a closed formula, given in Remark 6.6, allowed us to follow Terracini’s method in this setting.
Acknowledgments: I would like to thank my advisor, Dr. Alexandra Seceleanu, for her support and infinite patience in providing detailed commentary on earlier versions of this paper. Her help was especially instrumental in the final version of Appendix B and the proof of Lemma 6.9. I thank Benjamin Briggs for spending numerous hours helping me with the gory details of abstract tangent spaces, his help was crucial in my initial proof of Theorem 7.5. I thank Ola Sobieska for introducing me to [BR07] and Jack Jeffries for his help fine tuning my lattice point counting skills. I thank Mahrud Sayrafi for his help with some of my Macaulay2 code. I have benefited enormously from my conversations with Jack Huizenga, Nawaj KC, Claudiu Raicu, Andrew Soto Levins, and Ryan Watson. Finally, I thank Paolo Mantero and Tài Huy Hà for writing [HM21] and Giorgio Ottaviani and Maria Brambilla for writing [BO08], their papers are the main inspirations for this work.
2. Weighted projective space and the interpolation problem
2.1. Weighted projective varieties
Fix integers $a_{0},\ldots,a_{n}\in{\mathbb{N}}$.
Definition 2.1.
The weighted projective space ${\mathbb{P}}={\mathbb{P}}(a_{0},\ldots,a_{n})$ is a quotient of the affine space $\mathbb{A}_{k}^{n+1}$ under the equivalence relation
$(p_{0},\ldots,p_{n})\sim(\lambda^{a_{0}}p_{0},\ldots,\lambda^{a_{n}}p_{n})\text{ for }\lambda\in k\setminus\{0\}.$ (2.1)
Denoting $k^{\times}=k\setminus\{0\}$ the multiplicative group of $k$, and viewing the above as an action of $k^{\times}$ on $\mathbb{A}_{k}^{n+1}$ allows to write concisely
${\mathbb{P}}(a_{0},\ldots,a_{n})=(\mathbb{A}_{k}^{n+1}\setminus\{0\})/k^{\times}.$
The usual (or straight) projective space ${\mathbb{P}}_{k}^{n}$ can be recovered from this construction as ${\mathbb{P}}(1,1,\ldots,1)$. We reserve the notation ${\mathbb{P}}_{k}^{n}$ for straight projective space and ${\mathbb{P}}$ or ${\mathbb{P}}(a_{0},\ldots,a_{n})$ for weighted projective space.
Let $S=k[x_{0},\ldots,x_{n}]$ be a polynomial ring with (nonstandard) grading given by $\deg(x_{i})=a_{i}$. Observe that endowing $S$ with this grading is equivalent to giving an action on $\mathbb{A}_{k}^{n+1}$ as in Equation 2.1. Specifically, a polynomial $f\in S$ is homogeneous of degree $d$ if and only if
$f(\lambda^{a_{0}}x_{0},\ldots,\lambda^{a_{n}}x_{n})=\lambda^{d}f(x_{0},\ldots,x_{n})\text{ for all }\lambda\in k^{\times}.$ (2.2)
Therefore the condition $f(p)=0$ is well defined on equivalence classes of $\sim$ in Equation 2.1. One can thus alternatively define weighted projective space as the projective variety ${\mathbb{P}}=\operatorname{Proj}(S)$. More generally, if $I$ is a homogeneous ideal of $S$, then the projective variety $\operatorname{Proj}(S/I)$ is the quotient ${(V(I)\setminus\{0\})/k^{\times}}$. Throughout the paper $S$ always represents the (non-standard) graded homogenous polynomial ring of ${\mathbb{P}}(a_{0},\ldots,a_{n})$.
Much like the usual projective space, weighted projective space is covered by the affine charts $U_{i}=\{p\in{\mathbb{P}}\mid p_{i}\neq 0\}$. The coordinate ring of each $U_{i}$ is the ring of invariant polynomials $k[x_{0},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]^{\mu_{a_{i}}}$, where the group $\mu_{a_{i}}$ is identified with the subgroup of $k^{\times}$ of $a_{i}$-th roots of 1 and the group acts as follows: an element $\xi\in k$ with $\xi^{a_{i}}=1$ maps $x_{j}\mapsto\xi^{-a_{j}}x_{j}$. By Equation 2.2 a polynomial is invariant under this action if and only if its degree is divisible by $a_{i}$ leading to a ring isomomorphism
$k[x_{0},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}]^{\mu_{a_{i}}}\cong S[x_{i}^{-1}]_{0},f\mapsto\frac{f}{x_{i}^{\deg(f)/a_{i}}}.$ (2.3)
Thus $U_{i}$ is $\operatorname{Spec}(k[x_{0},\ldots,x_{i-1},x_{i+1},\ldots,x_{n}])/\mu_{a_{i}}=\mathbb{A}^{n}/\mu_{a_{i}}$.
Definition 2.2.
The $d$-th Veronese subring of $S$ is the graded ring
$S^{[d]}=\bigoplus_{i\geq 0}S_{id}.$ (2.4)
Although $S^{[d]}$ is a proper subring of $S$, $\operatorname{Proj}(S^{[d]})$ and $\operatorname{Proj}(S)$ are isomorphic projective varieties, because any homogeneous ratio in $S$ can be expressed as a homogeneous ratio in $S^{[d]}$. There are two different conventions in use about degrees in $S^{[d]}$: one can let the elements of the Veronese subring have the same degree they had in $S$, or one can divide degrees through by $d$. The advantage of the latter approach is that there exists a Veronese subring $S^{[D]}$ which is generated by its elements of the smallest degree ([Gro60, Lemma 2.1.6]); in this case, after dividing degrees by $D$, one can view $S^{[D]}$ as being standard graded. This gives rise to an embedding ${\mathbb{P}}\subset{\mathbb{P}}^{N}_{k}$, where $N=\dim_{k}(S_{D})-1$; see Remark B.2.
Another application of this ideal is a reduction to sets of weights that satisfy certain a simple condition which we now describe.
Definition 2.3.
A weighted projective space ${\mathbb{P}}(a_{0},\ldots,a_{n})$ is well formed if no $n-1$ of $a_{0},a_{1},\ldots,a_{n}$ have a common factor greater than $1$.
Remark 2.4.
Any weighted projective space is isomorphic to a well formed weighted projective space. Two weighted projective spaces with the same weights are isomorphic up to a reordering of the weights. Henceforth the space ${\mathbb{P}}(a_{0},\ldots,a_{n})$ is assumed to be well-formed with weights in increasing order $a_{0}\leq\cdots\leq a_{n}$ and $a_{i}\neq 1$ for some $0\leq i\leq n$.
2.2. Points in weighted projective space
In this section we describe the ideal defining one point in the weighted projective space and its properties.
Definition 2.5.
Let $p\in{\mathbb{P}}$ be a point. The defining ideal of $p$ is the ideal
$I_{p}=(f\in S\mid f\text{ homogeneous},f(p)=0).$
Lemma 2.6.
Fix a point $p\in{\mathbb{P}}$. Let $T=k[t]$ and consider the graded $k$-algebra homomorphism
$\varphi_{p}:S\to T,x_{i}\mapsto p_{i}t^{a_{i}}.$
Then $I_{p}=\ker(\varphi_{p})$ and $S/I_{p}\cong k[t^{a_{i}}\mid p_{i}\neq 0]$. Moreover, if $p,q\in{\mathbb{P}}$ are points such that for all $0\leq i\leq n$, $p_{i}=0$ if and only if $q_{i}=0$, then $I_{p}$ and $I_{q}$ are isomorphic $S$-modules.
Proof.
Since $\varphi_{p}$ is a degree-preserving homomorphism, $\ker(\varphi_{p})$ is a homogeneous ideal. If $f\in S$ is homogeneous of degree $d$ then $f\in\ker(\varphi_{p})$ if and only if $f(p_{i}t^{a_{i}})=0$ if and only if $t^{d}f(p)=0$ if and only if $f\in I_{p}$.
The map $\Theta:S\to S,\Theta(x_{i})=\frac{p_{i}}{q_{i}}x_{i}$ (where we make the convention $\frac{0}{0}=1$) is an automorphism on $S$ if and only if for all $0\leq i\leq n$, $p_{i}=0\iff q_{i}=0$. It satisfies $\Theta(I_{p})=I_{q}$ thus demonstrating an isomorphism $I_{p}\cong I_{q}$. ∎
In the case $p={\bf 1}=[1:1:\cdots:1]$ the ideal $I({\bf 1})=\ker(\varphi_{\bf 1})$ is called a monomial curve ideal. This class of ideals has been studied in [Her70], [Bre79] and [GKW91]. An expository survey of affine monomial curves and their computational aspects can be found in [Sen20]. The particular case of affine space curves is best understood. It was studied in [Her70].
Proposition 2.7 ([Her70, Den03]).
Consider positive integers $(a,b,c)$ with $\gcd(a,b,c)=1$. Let $r_{i}$ for $i=1,2,3$ be the smallest positive integer such that the following equations admit a solution in non-negative integers.
$\begin{split}r_{1}a&=k_{1}b+g_{1}c\\ r_{2}b&=k_{2}a+g_{2}c\\ r_{3}c&=k_{3}a+g_{3}b\end{split}$ (2.5)
Then the kernel of the $k$-algebra map
$\varphi:k[z,u,v]\to k[t],z\mapsto t^{a},u\mapsto t^{b},v\mapsto t^{c}$
is the ideal
$I=(z^{r_{1}}-u^{k_{1}}v^{g_{1}},u^{r_{2}}-z^{k_{2}}v^{g_{2}},v^{r_{2}}-z^{k_{3}}u^{g_{3}}).$
Moreover two of the binomials above agree up to sign if and only if $0\in\{k_{1},k_{2},k_{3},g_{1},g_{2},g_{3}\}$.
If in Equation 2.5 $k_{i}=0$ or $g_{i}=0$ for any $i$, then we say the weights $(a,b,c)$ satisfy the Herzog Criteria (HC).
Proposition 2.8.
Consider a well formed weighted projective plane $\mathbb{P}(a,b,c)$ with coordinate ring $S=k[z,u,v]$. Then the following hold:
  1. (1)
    If $p_{i}\neq 0$ for $1\leq i\leq 3$ then
    $p=[0:p_{1}:p_{2}]$$\Rightarrow$ $I_{p}=(z,p_{2}^{b}u^{c}-p_{1}^{c}v^{b})$
    $p^{\prime}=[p_{0}:0:p_{2}]$$\Rightarrow$ $I_{p^{\prime}}=(u,p_{2}^{a}z^{c}-p_{0}^{c}v^{a})$
    $p^{\prime\prime}=[p_{0}:p_{1}:0]$$\Rightarrow$ $I_{p^{\prime\prime}}=(v,p_{1}^{a}z^{b}-p_{0}^{b}u^{a}).$
  2. (2)
    Let $p=[p_{0}:p_{1}:p_{2}]$ with $p_{i}\neq 0$ for all $i$. Then
    $I_{p}=(p_{1}^{k_{1}}p_{2}^{g_{1}}z^{r_{1}}-p_{0}^{r_{1}}u^{k_{1}}v^{g_{1}},p_{0}^{k_{2}}p_{2}^{g_{2}}u^{r_{2}}-p_{1}^{r_{2}}z^{k_{2}}v^{g_{2}},p_{0}^{k_{3}}p_{1}^{g_{3}}v^{r_{3}}-p_{2}^{r_{3}}z^{k_{3}}u^{g_{3}})$
    where $r_{i}$, $k_{i}$ and $g_{i}$ are described in Equation 2.5
Proof.
Let $p=[0:p_{1}:p_{2}]$. The polynomial $p_{2}^{b}u^{c}-p_{1}^{c}v^{b}$ is irreducible because the only homogeneous polynomials in $k[u,v]$ of degree less than $bc$ are monomials. Observe that $J=(z,p_{2}^{b}u^{c}-p_{1}^{c}v^{b})\subseteq I_{p}$, and $J$ is a prime ideal of height 2. Since $I_{p}$ is also a prime ideal of height two, it follows that $J=I_{p}$. Other cases of $(1)$ are proven similarly.
Statement (2) follows from Lemma 2.6 and Proposition 2.7. Specifically, the proof of Lemma 2.6 yields that $I_{p}=\Theta(I_{\bf 1})$ where $\Theta(x_{i})=x_{i}/p_{i}$ and $I_{\bf 1}$ is described as $I$ in Proposition 2.7. ∎
Corollary 2.9.
All ideals defining points in the weighted projective plane ${\mathbb{P}}(a,b,c)$ are complete intersections if and only if $a,b,c$ satisfy (HC).
The next proposition shows that ideals defining points in certain higher dimensional weighted projective spaces are also complete intersections.
Proposition 2.10.
Consider ${\mathbb{P}}=\mathbb{P}(1,1,\ldots,1,a_{i+1},\ldots,a_{n})$ with $a_{j}\neq 1$ for all $i+1\leq j\leq n$ and let $S=k[x_{0},\ldots,x_{n}]$ be the coordinate ring of ${\mathbb{P}}$. Let $p=[p_{0}:\ldots:p_{n}]$ where $p_{t}\neq 0$ for some $0\leq t\leq i$. Then
$I_{p}=(p_{0}x_{t}^{a_{0}}-p_{t}^{a_{0}}x_{0},\ldots,p_{n}x_{t}^{a_{n}}-p_{t}^{a_{n}}x_{n}).$
Proof.
Note that $J=(p_{1}x_{t}^{a_{1}}-p_{t}^{a_{1}}x_{1},\ldots,p_{n}x_{t}^{a_{n}}-p_{t}^{a_{n}}x_{n})\subseteq I_{p}$ and $S/J\cong k[x_{t}]$ implies $J$ is a prime ideal. Further $\text{height}(J)=n$ because $J$ is a complete intersection with $n$ generators. Since $I_{p}$ is also a prime ideal of height $n$ the equality $J=I_{p}$ follows. ∎
We end this section by noting that finding explicit generators for the defining ideal of a point in ${\mathbb{P}}(a_{0},\ldots,a_{n})$ for $n\geq 3$ is much harder. In fact, outside of the case of Proposition 2.10 we do not know what the minimal generators of ideals of points are.
Problem 2.11.
Find explicit generators for the defining ideal of a general point in ${\mathbb{P}}(a_{0},\ldots,a_{n})$ for $n\geq 3$.
2.3. Hilbert function
Throughout ${\mathbb{N}}$ denotes the natural numbers (not including zero), ${\mathbb{N}}_{0}={\mathbb{N}}\cup\{0\}$, $k$ is a field and $S=k[x_{0},\ldots,x_{n}]$ is a polynomial ring in variables $x_{0},\ldots,x_{n}$ with coefficients in $k$. We endow $S$ with a grading given by $\deg(x_{i})=a_{i}\in{\mathbb{N}}$.
A graded $S$-module is a module that admits a decomposition $M=\bigoplus_{i\in{\mathbb{Z}}}M_{i}$, where $M_{i}$ denotes the $k$ vector space of homogeneous elements of $M$ of degree $i$. The Hilbert function of $M$ is the function $H_{-}(M):{\mathbb{N}}_{0}\to{\mathbb{N}}_{0}$ given by $H_{i}(M)=\dim_{k}M_{i}$.
Notation 2.12.
Throughout the paper we denote the Hilbert function of the polynomial ring $S$ by $s_{i}=\dim_{k}S_{i}$.
The integer $s_{d}$ is the number of monomials of degree $d$ in the polynomial ring $S$. Computing this number is equivalent to finding the number of non-negative integer solutions to the Diophantine equation
$\sum_{i=0}^{n}a_{i}t_{i}=d,\text{ where }t_{i}\in{\mathbb{N}}_{0}\text{ for all }0\leq i\leq n.$
The case of two variables has explicit formulas which are used repeatedly.
Example 2.13.
Let $S=k[z,u]$ with $\deg(z)=a$ and $\deg(y)=b$ where $\gcd(a,b)=1$. Let $p$ and $q$ be positive integers such that $aq-bp=1$. By [BR07, Corollary 1.6] we have
$s_{d}=\begin{cases}\left\lfloor\frac{qn}{b}\right\rfloor-\left\lfloor\frac{pn}{a}\right\rfloor&\text{ if }a\nmid d\\ \left\lfloor\frac{qn}{b}\right\rfloor-\frac{pn}{a}+1&\text{ if }a\mid d.\end{cases}$ (2.6)
In particular for $a=1$ this yields
$s_{d}=\left\lfloor\frac{d}{b}\right\rfloor+1.$ (2.7)
Remark 2.14.
For a graded $S$-module $M$, where $x_{i}$ is a non-zero divisor of degree $a_{i}$ on $M$ we have the equation
$H_{d}(M)=H_{d-a_{i}}(M)+H_{d}(M/(x_{i}\cdot M)).$ (2.8)
Equation 2.8 is extensively used in the case of ${M=S/I_{X}}$ where $I_{X}$ defines a set of points contained in $U_{i}$.
2.4. The interpolation problem
Based on our assumptions on the field, we can define the symbolic powers of a point as follows.
Definition 2.15.
Let $p\in{\mathbb{P}}$ be a point. The ideal defining the multiple point $mp$ (also termed a fat point ideal or a symbolic power ideal) is
$I_{p}^{(m)}=\left\{f\in S\mid\frac{\partial^{m-1}f}{\partial x_{0}^{u_{0}}\cdots\partial x_{n}^{u_{n}}}(p)=0\text{ where }\sum u_{t}=m-1\right\}.$
Note when the defining ideal of a simple point is a complete intersection the equality $I_{p}^{(m)}=I_{p}^{m}$ holds for all $m\in{\mathbb{N}}$. Moreover, for a set $X=\{m_{i}p_{i}\}_{i=1}^{r}$ of multiple points the defining ideal is
$I_{X}=\bigcap_{i=1}^{r}I_{p_{i}}^{(m_{i})}.$
The ideal above contains the equations of hypersurfaces vanishing to order at least $m_{i}$ at each point $p_{i}$. This interpretation leads to the following.
Problem 2.16.
(Interpolation Problem) Find the dimension of the vector space of hypersurfaces of a degree $d$ vanishing to order at least $m_{i}$ at each points $p_{i}$ in a given finite set. Equivalently, find the Hilbert function the coordinate ring of a set $X=\{m_{i}p_{i}\}_{i=1}^{r}$ of multiple points.
Using Definition 2.15 for a set $X=\{m_{i}p_{i}\}_{i=1}^{r}$ of points with mixed multiplicity we notice the quantity $\sum_{i=1}^{r}\binom{n+m_{i}-1}{n}$ is the number of conditions imposed on degree $d$ hypersurfaces by the vanishing of all partial derivatives of appropriate order. It is natural to inquire if these conditions are linearly independent. Clearly, the vanishing conditions can not be linearly independent if their number exceeds $s_{d}$, the dimension of the space of all hypersurfaces of degree $d$. This brings about the next definition which signifies the expected number of linearly independent conditions imposed by $X$.
Definition 2.17 ($\text{AH}_{n}(d)$ property).
Let $X=\{m_{i}p_{i}\}_{i=1}^{r}\subseteq{\mathbb{P}}(a_{0},\ldots,a_{n})$ be a set of $r$ points with defining ideal $I_{X}\subseteq S$. For $d\in{\mathbb{N}}$, we say $X$ has property $\text{AH}_{n}(d)$ or $X$ is $\text{AH}_{n}(d)$ if
$H_{d}(R/I_{X})=\min\left\{s_{d},\sum_{i=1}^{r}\binom{n+m_{i}-1}{n}\right\}.$
Proposition 2.19 states a well-known equivalent formulations of interpolation in terms of the evaluation map which is now introduced.
Definition 2.18 (Evaluation Map).
Let $X=\{m_{i}p_{i}\}_{i=1}^{r}\subseteq{\mathbb{P}}(a_{0},\ldots,a_{n})$ be a set of $r$ double points with $m=\sum_{i=1}^{r}\binom{n+m_{i}-1}{n}$. For any $d\in{\mathbb{N}}$, the evaluation map on $X$ in degree $d$ is the map
$\operatorname{ev}_{X}:$ $(S)_{d}$ $\to k^{m}$
$f$ $\to\left(\frac{\partial^{m_{i}-1}f}{\partial x_{0}^{u_{0}}\cdots\partial x_{n}^{u_{n}}}(p_{i})=0\mid\sum u_{t}=m_{i}-1\right)_{i=1}^{r}.$
Proposition 2.19.
For $X\subseteq{\mathbb{P}}(a_{0},a_{1},\ldots,a_{n})$ as defined in Definition 2.18 and any $d\in{\mathbb{N}}_{0}$ the following are equivalent:
  1. (1)
    $X$ is $\text{AH}_{n}(d)$,
  2. (2)
    $\operatorname{ev}_{X}:(S)_{d}\to k^{m}$ has full rank.
Remark 2.20.
Since $I_{X}$ is the kernel of the evaluation map and $S/I_{X}$ is its image, it is clear that
$H_{d}(S/I_{X})\leq\min\{s_{d},m\}.$
Therefore if $H_{d}(S/I_{X})$ is equal to $s_{d}$ or $m$ then that value is automatically the minimum of the two.
3. Mixed point interpolation in the weighted projective line
This subsection focuses on interpolation in ${\mathbb{P}}(a,b)$, where $\gcd(a,b)=1$ and coordinate ring $S=k[z,u]$ where $\deg(z)=a$ and $\deg(u)=b$.
Proposition 3.1.
Let $I$ be the defining ideal of a point $p=[p_{0}:p_{1}]$. Then
$I_{p}=\begin{cases}(p_{1}^{a}z^{b}-p_{0}^{b}u^{a})&\text{if }p_{0},p_{1}\neq 0\\ (z)&\text{if }p_{0}=0\text{ and }p_{1}\neq 0\\ (u)&\text{if }p_{0}\neq 0\text{ and }p_{1}=0.\end{cases}$
Proof.
Note $I_{p}$ is an unmixed height $1$ prime ideal, which implies $I_{p}$ is principal. Therefore it suffices to find an element of smallest degree in $I_{p}$. If $p_{0}=0$ or $p_{1}=0$ this element is evidently $z$ and $u$, respectively. Assuming $p_{0},p_{1}\neq 0$ implies no pure power of $z$ or $u$ belongs to $I_{X}$. Because $\gcd(a,b)=1$ and $(p_{1}^{a}z^{b}-p_{0}^{b}u^{a})\subseteq I_{X}$, the least degree homogenous element in $I_{X}$ is $p_{1}^{a}z^{b}-p_{0}^{b}u^{a}$. ∎
Proposition 3.2.
A set of general points with mixed multiplicity, $X=\{r_{i}p_{i}\}_{i=1}^{n}\subseteq{\mathbb{P}}(a,b)$, is $\text{AH}_{1}(d)$ for all $d\in{\mathbb{N}}$, that is
$H_{d}(S/I_{X})=\min\{s_{d},\sum_{i=1}^{n}r_{i}\}.$ (3.1)
Proof.
It suffices to demonstrate one such set. (See Remark 5.1.) Let $X=\{r_{1}p_{1},\ldots,r_{n}p_{n}\}$ where $p_{i}=[p_{i}^{(0)}:p_{i}^{(1)}]$ and $[0:p_{1}]\not\in X$. Recall that $s_{d}=H_{d}(k[z,u])$ is the number of solutions to
$ax+by=d\text{ where }x,y\in{\mathbb{N}}_{0}.$ (3.2)
We wish to interpret $H_{d}(S/I_{X})$ in terms of Equation 3.2. For any $i$, Proposition 3.1 implies $I_{p_{i}}=((p_{i}^{(0)})^{b}z^{b}-((p_{i}^{(1)})^{a}u^{a})$ is a principal prime ideal and thus
$I_{X}$ $=$ $\bigcap_{i=1}^{n}I_{p_{i}}^{(r_{i})}=\bigcap_{i=1}^{n}\left((p_{i}^{(0)})^{b}z^{b}-((p_{i}^{(1)})^{a}u^{a})^{(r_{i})}\right)$
${*}$ $\left(\prod_{i=1}^{n}((p_{i}^{(0)})^{b}z^{b}-((p_{i}^{(1)})^{a}u^{a})^{r_{i}}\right).$
Identity * follows since each ideal is a complete intersection. Replacing $I_{X}$ by its initial ideal $(u^{ar})$, where $r=\sum_{i=1}^{n}r_{i}$, we notice that
$S/I_{X}=\text{span}_{k}\{z^{x}u^{y}\;|\;0\leq y\leq ar-1,\;0\leq x\}\underset{k-\text{vector space}}{\cong}k[z]\otimes_{k}k[u]/(u^{ar}).$
Hence $H_{d}(S/I_{X})$ is the number of solutions to Equation 3.2 with the restriction that $10\leq y\leq ar-$ 1. We will show
$H_{d}(S/I_{X})=\begin{cases}s_{d}&\text{ if }d<b(ar-1)\\ r&\text{ if }d\geq b(ar-1).\end{cases}$ (3.3)
Suppose $d<b(ar-1)$. Then $20\leq y\leq ar-$ 2 regardless of whether we are computing $s_{d}$ or $H_{d}(S/I_{X})$, which implies that $H_{d}(S/I_{X})=s_{d}$ holds in this case. Let $d\geq b(ar-1)$. Then $y=ar-i$ for $1\leq i\leq ar$. It suffices to count how many values of $i$ yield a solution. From Equation 3.2 we get
$ax+b(ar-i)=d\Rightarrow bi+d\equiv 0\pmod{a}\Rightarrow(-b)^{-1}d\equiv i\pmod{a}$
Hence $[i]_{a}=[(-b)^{-1}d]_{a}$ and $1\leq i\leq ar$. Therefore $H_{d}(S/I_{X})$ is equal to the number of distinct coset representatives of $[(-b)^{-1}(d)]_{a}$ in the interval $[1,ar]$; of which there are $r$. Hence $H_{d}(S/I_{X})=r$ for $d\geq b(ar-1)$. Combining Remark 2.20 and Equation 3.3 implies Equation 3.1
4. Simple point interpolation in the weighted projective space
The objective of this section is to establish simple point interpolation in ${\mathbb{P}}(a_{0},\ldots,a_{n})$. Theorem 4.1 is implicitly used in subsequent proofs of Proposition 5.7, Theorem 6.1 and Theorem 7.7. Our proof is inspired by the one in [Hui18].
Theorem 4.1.
Let ${\mathbb{P}}={\mathbb{P}}(a_{0},a_{1},\ldots,a_{n})$ and $X=\{p_{i}\}_{i=1}^{r}$ be a set of general simple points. $X$ is $\text{AH}_{n}(d)$ for all $d$, that is
$H_{d}(S/I_{X})=\min\{s_{d},r\}$ (4.1)
Proof.
Let $k$ be an algebraically closed field. We use induction on $r$, the number of points in $X$. Let $r=1$ and choose $p=[p^{(0)}:\cdots:p^{(n)}]$ such that $p^{(j)}\neq 0$ for all $j$. By Lemma 2.6, $S/I_{p}\cong k[t^{a_{j}}\mid 0\leq j\leq n]$. The observations below implies Equation 4.1 for $X=\{p\}$ where $A=\{a_{j}\}_{j=0}^{n}$ and $\mathcal{N}(A)$ denotes the numerical semigroup generated $A$.
$H_{d}(S/I_{X})=\begin{cases}1&\text{ if }d\in\mathcal{N}(A)\\ 0&\text{else}\end{cases},\;\;s_{d}=\begin{cases}\geq 1&\text{ if }d\in\mathcal{N}(A)\\ 0&\text{else}\end{cases}$
Let $X^{\prime}$ be a set of size $r-1$ for which Equation 4.1 holds for all $d$. We claim for any fixed $D$ and some choice of a point $p$, there exists a set $X=X^{\prime}\cup\{p\}$ such that Equation 4.1 holds for $H_{D}(S/I_{X})$. If $s_{D}\leq r-1$, choose $p\not\in X^{\prime}$. Then
$(I_{X})_{D}\subseteq(I_{X^{\prime}})_{D}=0\Rightarrow H_{D}(S/I_{X})=s_{D}<r.$
If $r-1<s_{D}$ then there exists $f\in(I_{X^{\prime}})_{D}$ such that $f\neq 0$. Choose $p\in{\mathbb{P}}$ such that $p\not\in X^{\prime}\cup V(f)$. Such choice is possible because $X\cup V(f)\neq{\mathbb{P}}$ by the Nullstellensatz. In particular, $f(p)\neq 0$. Observe that $(I_{X})_{D}=\{g\in(I_{X^{\prime}})_{D}\;|\;g(p)=0\}$ is the kernel of the map
$(I_{X^{\prime}})_{D}\xrightarrow{}k\text{ where }g\mapsto g(p),$
which is surjective since $f(p)\neq 0$. Hence $H_{D}(I_{X})=H_{D}(I_{X^{\prime}})-1$ and
$H_{D}(S/I_{X})=H_{D}(S)-H_{D}(I_{X})=H_{D}(S)-H_{D}(I_{X^{\prime}})+1=H_{D}(S/I_{X^{\prime}})+1=r,$
where $r=\min\{s_{D},r\}$ by choice of $D$. This proves the claim.
Because the construction of the set $X$ depended on the degree $D$, the claim above shows for a fixed D there exists a set of size $r$ that satisfies Equation 4.1. Equivalently, there exists a non-empty open set in the parameter space ${\mathbb{P}}(a_{0},\ldots,a_{n})^{r}$, denoted $C_{D}$, whose points satisfy Equation 4.1. Next we show the existence of a single set $Y$ of size $r$, independent of $D$, that satisfies Equation 4.1 for all degrees $d$.
We argue to obtain an integer $t$ such that $r\leq s_{d}$ for all $d\geq t-a_{n}$. The well-formedness of $a_{i}$’s, Remark 2.4, implies $\gcd(a_{0},a_{1},\ldots,a_{n})=1$. Since $r$ is fixed and the Hilbert function of $S$ for $d\gg 0$ is given by a quasi-polynomial
$H_{d}(S)=P_{i}(d)\text{ for }d\equiv i\pmod{\operatorname{lcm}(a_{0},\ldots,a_{n})}$
where each $P_{i}$ is a polynomial of degree $n\geq 1$ such that $P_{i}(d)>0$ for $d\gg 0$, it follows that each sequence $\{P_{i}(d)\}_{d}$ is eventually increasing. Thus the existence of the desired $t$ is established. Moreover choose the smallest such integer $t$. Consider
$C=\underset{0\leq D\leq t+\cdot a_{n}}{\bigcap}C_{D}\cap\left({\mathbb{P}}(a_{0},\ldots,a_{n})^{r}\setminus Z\right),$
where $Z$ represents the closed set of $r$-tuples of points in ${\mathbb{P}}(a_{0},\ldots,a_{n})$ where each coordinate hyperplane $V(x_{j})$ contains at least one of the points or where any two or more of the points coincide.
By construction $C$ is a non-empty open set since it is the intersection of finitely many non-empty open sets and points of $C$ satisfy Equation 4.1 for all $0\leq D\leq t+a_{n}$. We finish the proof by showing this implies points of $C$ satisfy Equation 4.1 for all $d$. Fix an $r$-tuple of $C$ which corresponds to a set $Y\subseteq{\mathbb{P}}$ of $r$ points. Notice some variable $x_{j}$ is a non-zero divisor on $S/I_{Y}$ and Equation 2.8 is now applicable. Next we show
$H_{d}(S/(I_{Y}+x_{j}))=0\text{ for }t\leq d\leq t+a_{n}.$ (4.2)
Let $d=t+\gamma$ where $0\leq\gamma\leq a_{n}$. Note $t-a_{n}\leq t-a_{j}$ and $H_{d^{\prime}}(S/I_{Y})=r$ for all $t-a_{n}\leq d^{\prime}\leq t+a_{n}$ by choice of $t$. Plugging $d$ in Equation 2.8 gives
$H_{t+\gamma}(S/I_{Y})=H_{t+\gamma-a_{j}}(S/I_{Y})+H_{t+\gamma}(S/(I_{Y}+x_{j})).$
Which implies Equation 4.2. Finally, $S/(I_{Y}+x_{n})$ is a graded cyclic $S$-module and since $[S/(I_{Y}+x_{n})]_{d}\subseteq[S/(I_{Y}+x_{n})]_{d-a_{n}}S$ for all $d\geq a_{n}$, Equation 4.2 yields
$H_{d}(S/(I_{Y}+x_{j}))=0\text{ for }t+a_{n}\leq d.$
For an arbitrary choice of $Y$ we have shown
$H_{d}(S/I_{Y})=\begin{cases}r&d\geq t-a_{n}\\ s_{d}&d<t-a_{n}\end{cases}\overset{*}{=}\min\{r,s_{d}\}\text{ for all }d.$ (4.3)
Where * follows from Remark 2.20. Thus a set of $r$ general simple points satisfies Equation 4.1. ∎
5. Double point interpolation in the weighted projective plane
This section focuses on interpolation in weighted projective planes of the form ${\mathbb{P}}(1,b,c)$, with coordinate ring $k[z,u,v]$ with $\deg(z)=1$, $\deg(u)=b$ and $\deg(v)=c$. We work with the evaluation map to show general double points have the $\text{AH}_{2}(d)$ property for all $d$, which recall is to show
$H_{d}(S/I_{X})=\min\{s_{d},3\cdot|X|\}.$
Remark 5.1.
It is well-known that in order to prove a general set of points has the $\text{AH}_{n}(d)$ property, it suffices to demonstrate one such set. One can further limit the size of this set to at most two values. For a fixed $n$ and $d$, it suffices to prove the property $\text{AH}_{n}(d)$ for one set of double points $X$, where $|X|=r$ and $r=\left\lfloor\frac{1}{n+1}s_{d}\right\rfloor$ and/or $r=\left\lceil\frac{1}{n+1}s_{d}\right\rceil$; see [HM21, Corrollary D.4-D.5]. We employ this in the weighted projective plane where $n=2$.
Remark 5.2.
The existence of a weight of $1$ ensures the following:
  1. (1)
    The defining ideal of a point is a complete intersection. (Corollary 2.9.)
  2. (2)
    We can choose $U_{0}={\mathbb{P}}(1,b,c)\setminus V(z)$ as our locus of interpolation. Note $U_{0}$ is a non-empty open dense set in the parameter space, in this case ${\mathbb{P}}(1,b,c)$. Hence our notion of general refers to sets of points that are contained in $U_{0}$ and is made more precise in different results. The advantage is if $X\subseteq U_{0}$ is a set of points and $H_{d}(S/I_{X})=(n+1)r$ for some $d$, then Definition 2.18 and Equation 2.8 imply $H_{d^{\prime}}(S/I_{X})=(n+1)r$ for all $d^{\prime}\geq d$.
  3. (3)
    The coordinate ring of ${\mathbb{P}}(1,b,c)$, $S=k[z,u,v]$, has non-decreasing Hilbert function. Since $z$ is a degree $1$ non-zero divisor, Equation 2.8 yields $H_{d}(S)\geq H_{d-i}(S)$ for all $i$ hence $\{s_{i}\}$ forms a non-decreasing sequence.
Proposition 5.4 shows to find a space in which $r$ general double points are $\text{AH}_{2}(d)$ for all $d$, it is necessary for $c$ to not be too large compared to $b$ and $r$. Lemma 5.3 provides a numerical observation towards this end.
Lemma 5.3.
Fix positive integers $b,c$ and $r$ such that $rb<c<(r+1)b$. For the coordinate ring $S$ of ${\mathbb{P}}(1,b,c)$, the smallest $d$ such that $s_{d}=3r$ is $d=(r-1)b+c$.
Proof.
For $d=(r-1)b+c$, observe that $\deg(v^{2})>d$ and we get
$k[z,u,v]_{d}=k[z,u]_{d}\oplus v\cdot k[z,u]_{d-c}.$ (5.1)
This implies by Equation 2.7 the claimed identity
$s_{d}=\left\lfloor\frac{d}{b}\right\rfloor+1+\left\lfloor\frac{d-c}{b}\right\rfloor+1=3r.$
To see the minimality of $d$, note for $d^{\prime}<d$, an analogous decomposition to Equation 5.1 implies
$s_{d^{\prime}}=\left\lfloor\frac{d^{\prime}}{b}\right\rfloor+1+\left\lfloor\frac{d^{\prime}-c}{b}\right\rfloor+1<3r.\qed$
Proposition 5.4.
If $r$ general double points are $\text{AH}_{2}(d)$ for all $d$ in ${\mathbb{P}}(1,b,c)$, then
$c$$\leq$ $(r+1)\text{ for }b=1$
$c$ $<$ $(r+1)b\text{ for }b\neq 1$
Proof.
Let $X=\{2p_{i}\}_{i=1}^{r}$ be a set of points such that $p_{i}\neq[0:1:0]$ Assume for contradiction that $c>(r+1)b$. It suffices to show $s_{2rb}\leq 3r$. This claim implies $(I_{X})_{2rb}=0$. But by Proposition 2.8, the square of the ideal of each point, $I_{p_{i}}^{2}$, contains a generator, $f_{i}$, of degree $2b$. Clearly $F=\prod_{i=1}^{r}f_{i}\in I_{X}$ and $\deg(F)=2rb$. This contradicts $X$ being $\text{AH}_{2}(2rb)$.
To verify $s_{2rb}\leq 3r$, observe $\deg(v^{2})=2c>2rb$ and Equation 5.1 yields
$s_{2rb}=\left\lfloor\frac{2rb}{b}\right\rfloor+1+\left\lfloor\frac{2rb-c}{b}\right\rfloor+1\leq 2r+1+(r-2)+1\leq 3r.\qed$
5.1. Small number of points in ${\mathbb{P}}(1,b,c)$
Remark 5.5 explains our proof strategy for the rest of Section 5.
Remark 5.5.
Let $M_{d}$ denote the matrix of $\operatorname{ev}_{X}:(S)_{d}\to k^{3r}$. We wish to demonstrate one set of double points $X$, of size $r$, that is $\text{AH}_{2}(d)$ for all $d$. If there is a degree $d$ such that $s_{d}=3r$, then it suffices to check the rank of $M_{d}$ is maximal for smallest such $d$. By Remark 5.2 and Proposition 2.19, it follows that $X$ is $\text{AH}_{2}(d^{\prime})$ for all $d^{\prime}\geq d$. Moreover, for $d^{\prime\prime}\leq d$ having a weight of $1$ implies
$(I_{X})_{d}=0\Rightarrow(I_{X})_{d^{\prime\prime}}=0.$
Combining this with Remark 5.2 – part 3 implies $X$ is $\text{AH}_{2}(d^{\prime\prime})$ for $d^{\prime\prime}\leq d$. For certain numbers $r$ the degree where $s_{d}=3r$ is identified in Lemma 5.3.
We show $M_{d}$ has full rank by exhibiting a set of points at which $\det(M_{d})$, a polynomial in $2r$ variables corresponding to the coordinates of $r$ points, is not zero. Thus $\det(M_{d})$ is not the zero polynomial. Hence the notion of general will refer to points whose first coordinate is nonzero and last two coordinates are outside of $V(\det(M_{d}))\subseteq{\mathbb{P}}^{2r}$.
In Proposition 5.6 and Proposition 5.7 we give necessary and sufficient condition for $1$ point and $2$ points to be $\text{AH}_{2}(d)$ for all $d$.
Proposition 5.6.
One general double point in ${\mathbb{P}}(1,b,c)$ is $\text{AH}_{2}(d)$ for all $d$ if and only if
  1. (1)
    $c=2$ when $b=1$,
  2. (2)
    $b<c<2b$ when $b\neq 1$.
Proof.
()($\Rightarrow$)( ⇒ ) This is a direct consequence of Remark 2.4 and Proposition 5.4.
()($\Leftarrow$)( ⇐ ) Let $X=\{2p\}$ with $p\not\in V(z)$. When $b=1$ and $c=2$ observe that $H_{2}(S/I_{X})=3$ because $s_{2}=4$ and $H_{2}(I_{X})=1$. Further $H_{1}(S/I_{X})=1$. By a similar reasoning to Remark 5.5 $X$ is $\text{AH}_{2}(d)$ for all $d$. Suppose $b<c<2b$. By Lemma 5.3 and Remark 5.5, it suffices to check $\text{AH}_{2}(d)$ for $X$ in degree $d=c$. By Proposition 2.8, the smallest degree of a generator in $I_{X}$ is $2b>c$. Hence $H_{c}(S/I_{X})=3=\min\{s_{c},3\}$. ∎
Our results so far have been characteristic free. In Proposition 5.7 to ensure certain determinants are non-zero we make assumptions on $\operatorname{char}(k)$.
Proposition 5.7.
Let $k$ be an infinite field with $\operatorname{char}(k)\not\in\{2,3,b,b+c\}$. Two general double points in ${\mathbb{P}}(1,b,c)$, where $(1,b,c)\neq(1,2,3)$, are $\text{AH}_{2}(d)$ for all $d$ if and only if
  1. (1)
    $b=1$ and $c=2$ or $c=3$,
  2. (2)
    $2b<c<3b$, or
  3. (3)
    $\frac{3}{2}b<c<2b$
Proof.
()($\Leftarrow$)( ⇐ ) If $b=1$, $c=3$ and $X$ is a set of two double points, one can use the defining equations of the points to show $H_{3}(S/I_{X})=5=s_{3}$ and $H_{4}(S/I_{X})=6$. By a similar reasoning to Remark 5.5, $X$ is $\text{AH}_{2}(d)$ for all $d$. Next let $2b<c<3b$ or b=1 and c=2. Lemma 5.3 and direct computations imply $s_{b+c}=6$ and this is the smallest degree in which $S$ has $6$ monomials: $z^{b+c},z^{c}u,z^{c-b}u^{2},z^{c-2b}u^{3},z^{b}v,uv$. The evaluation matrix with respect to this basis for points with coordinates $[1:p_{1}:p_{2}]$ and $[1:q_{1}:q_{2}]$ is
$M_{b+c}=\left[\begin{matrix}b+c&cp_{1}&(c-b)(p_{1})^{2}&(c-2b)\,(p_{1})^{3}&b\,p_{2}&0\\ 0&1&2\,p_{1}&3\,(p_{1})^{2}&0&p_{2}\\ b+c&cq_{1}&(c-b)(q_{1})^{2}&(c-2b)\,(q_{1})^{3}&b\,q_{2}&0\\ 0&1&2\,q_{1}&3\,(q_{1})^{2}&0&q_{2}\\ 0&0&0&0&1&p_{1}\\ 0&0&0&0&1&q_{1}\\ \end{matrix}\right]$
with determinant $\det(M_{b+c})=(-1)(p_{1}-q_{1})^{5}(b+c)^{2}\neq 0$.
Finally suppose $\frac{3}{2}b<c<2b$. Adapting the proof of Lemma 5.3, one can show $d=3b$ is the first degree where $s_{d}=6$. The evaluation matrix with respect to the basis $z^{3b},z^{2b}u,z^{b}u^{2},u^{3},z^{2b-c}uv,z^{3b-c}v$ and the points fixed before is
$M_{{3b}}=\left[\begin{matrix}3\,b&2\,b\,p_{1}&b\,(p_{1})^{2}&(p_{1})^{3}&(2b-c)\,p_{1}\,p_{2}&(3b-c)\,p_{2}\\ 0&1&2\,p_{1}&3\,(p_{1})^{2}&p_{2}&0\\ 0&0&0&0&p_{1}&1\\ 3\,b&2\,b\,q_{1}&b\,(q_{1})^{2}&(q_{1})^{3}&(2b-c)\,q_{1}\,q_{2}&(3b-c)\,q_{2}\\ 0&1&2\,q_{1}&3\,(q_{1})^{2}&q_{2}&0\\ 0&0&0&0&q_{1}&1\\ \end{matrix}\right]$
where $\det(M_{3b})=3b(p_{1}-q_{1})^{3}\left[3\,b\,(p_{1}-q_{1})^{2}-2\,p_{1}^{2}-2\,p_{1}\,q_{1}-2\,q_{1}^{2}\right]$.
()($\Rightarrow$)( ⇒ ) Suppose a set of two general points, $X$, is $\text{AH}_{2}(d)$. General in this context refers to $X$ not containing any of the points $[0:1:0]$ or $[1:0:0]$. When $b$ is $1$ the result is immediate from Proposition 5.4 so suppose $b\neq 1$ Proposition 5.4 results in $c<3b$ and if $2b<c<3b$ we are done. Suppose $b<c<2b$. We wish to show $\frac{3}{2}b<c$. The equality $3b=2c$ is impossible by well-formedness and the hypothesis which excludes the possibility $b=2,c=3$. Therefore one may assume for contradiction that $2c<3b$. Direct calculations show $s_{2c}=6$. By assumption
$H_{2c}(S/I_{X})=\min\{6,s_{2c}\}=6.$
This implies $(I_{X})_{2c}=0$. We arrive at a contradiction by constructing a non-zero curve in $(I_{X})_{2c}$. Since $b<c<2b$ it follows $s_{c}=3$ and thus there is a degree $c$ curve with defining equation $F$ passing through the two points. Since $F^{2}\in[I_{X}]_{2c}$ we have reached the desired contradiction. ∎
5.2. A sufficient condition for $r$ points
The main goal of this subsection is to give a sufficient condition for the weights so that $r$ general double points in ${\mathbb{P}}(1,b,c)$ are $\text{AH}_{2}(d)$ for all $d$. The proof of Proposition 5.8 is a generalization of case 1 in Proposition 5.7.
Proposition 5.8.
Let $r\in{\mathbb{N}}$. If $rb<c<(r+1)b$ then a set of $r$ general double points in ${\mathbb{P}}(1,b,c)$ is $\text{AH}_{2}(d)$ for all $d$.
Proof.
Let $X=\{2p_{i}\}$, where the points $p_{i}=[1:p_{i}^{(1)}:p_{i}^{(2)}]$ for $1\leq i\leq r$ satisfy the following conditions:
  1. (1)
    $p_{i}^{(1)}\neq 0$ and $p_{i}^{(2)}\neq 0\text{ for all }i$,
  2. (2)
    $p_{i}^{(1)}\neq p_{i^{\prime}}^{(1)}$ for all $i\neq i^{\prime}$.
Let $c=rb+j$ and $d=(r-1)b+c=(2r-1)b+j$ which, by Lemma 5.3, is the smallest degree $d$ where $s_{d}=3r$. By Remark 5.5 it suffices to show $H_{d}(S/I_{X})=3r$.
We order the monomials of degree $d$ as follows
$z^{((2r-1)-s)b+j}u^{s},z^{((r-1)-t)b}u^{t}v\text{ where }0\leq s\leq 2r-1,\;\;0\leq t\leq r-1.$
The rows of the evaluation map correspond to $\frac{\partial}{\partial z}|_{p_{i}},\frac{\partial}{\partial u}|_{p_{i}}$ and $\frac{\partial}{\partial v}|_{p_{i}}$ and the columns to monomials as ordered. Based on our ordering of the monomials, the first $2r$ monomials do not contain $v$ and the last $r$ do.
The matrix $M_{d}$ of the evaluation map with rows indexed by $\frac{\partial}{\partial z}|_{p_{i}},\frac{\partial}{\partial u}|_{p_{i}}$ in the first block and $\frac{\partial}{\partial v}|_{p_{i}}$ in the second block becomes
$\left[\begin{array}[]{c|c}\begin{array}[]{cccc}(2k-1)b+j&((2k-2)b+j)p_{1}^{(1)}&\cdots&(p_{1}^{(1)})^{2k-1}\\ \vdots&\vdots&&\vdots\\ (2k-1)b+j&((2k-2)b+j)p_{k}^{(1)}&\cdots&j(p_{k}^{(1)})^{2k-1}\end{array}&**\\ \hline\cr\mbox{\Large 0}&\begin{array}[]{ccccc}1&p_{1}^{(1)}&(p_{1}^{(1)})^{2}&\ldots&(p_{1}^{(1)})^{k-1}\\ \vdots&\vdots&\vdots&&\vdots\\ 1&p_{k}^{(1)}&(p_{k}^{(1)})^{2}&\ldots&(p_{k}^{(1)})^{k-1}\end{array}\end{array}\right]$
The $2r\times 2r$ minor in the upper left quadrant corresponds to the evaluation map in degree $d$ for the points $[1:p_{i}^{(1)}]\in\mathbb{P}(1,b)$. Assumption $(1)$ allows us to view the points in ${\mathbb{P}}(1,b)$ and this minor is non-zero by Proposition 3.2. The bottom right block is the $r\times r$ Vandermonde minor which is non-zero by assumption $(2)$. Hence $M_{d}$ has full rank which completes the proof. ∎
6. The Alexander–Hirschowitz theorem in ${\mathbb{P}}(1,2,3)$
This section is dedicated to the proof of the following result.
Theorem 6.1 (Alexander Hirschowitz theorem in ${\mathbb{P}}(1,2,3)$).
Let $k$ be a field where $\operatorname{char}(k)\not\in\{2,3,5\}$. Let $X\subseteq{\mathbb{P}}(1,2,3)$ be a set of $r$ general double points with defining ideal $I_{X}$. Then $X$ is $\text{AH}_{2}(d)$ for all $d$ with no exceptions. That is
$H_{d}(S/I_{X})=\min\{s_{d},3r\}\text{ for all }d.$
Proof.
We use strong induction on $d$. By Remark 5.1, for each $d$ and for each of the values $r=\left\lfloor\frac{1}{3}s_{d}\right\rfloor$ and $r=\left\lceil\frac{1}{3}s_{d}\right\rceil$ it suffices to demonstrate one set of size $r$ that is $\text{AH}_{2}(d)$. Proposition 6.4 proves $6$ base cases for the induction, that is any finite set of general double point is $\text{AH}_{2}(d)$ for $0\leq d\leq 5$.
For any $d\geq 6$, and either value of $r$, we satisfy the hypothesis of Theorem 6.2 which constructs a set of size $r$ that is $\text{AH}_{2}(d)$. This entails showing:
  • The numerical condition of Theorem 6.2, Equation 6.1, is satisfied by Lemma 6.8. Let $q$ denote the positive integer mentioned in this condition.
  • Part (1) of Theorem 6.2 is the content of Proposition 3.2 when $r_{i}=2$ for all $i$.
  • Part (2) of Theorem 6.2 follows because by the induction hypothesis a set of general double points of any size in ${\mathbb{P}}(1,2,3)$ is $\text{AH}_{2}(t)$ for $11\leq t\leq d-$ 1. In particular, a set of $r-q$ general double points in ${\mathbb{P}}(1,2,3)$ is $\text{AH}_{2}(d-i)$ and $\text{AH}_{2}(d-2i)$ for $i\in\{1,2,3\}$. In Lemma 6.9 we show, using Lemma 6.3 and the previous considerations, that adding $q$ general simple points lying on a projective line, to the set of $r-q$ general double point preserves the $\text{AH}_{2}(d-i)$ property. This constructs the desired set in part (2) of Theorem 6.2.
Theorem 6.2 is the main inductive step used in Theorem 6.1 and in the standard graded case is proven in [HM21] and [BO08]. The assumption on the ambient space in Theorem 6.2 allows us to conclude the defining ideal of a simple point is a complete intersection (Corollary 2.9, Proposition 2.10), a fact which is fundamental to the proof presented in [HM21]. In particular when $n\geq 3$ all general points in ${\mathbb{P}}$ are considered in $U=\bigcup_{\{j\;\mid\;a_{j}=1\}}U_{j}$ to ensure simple points are complete intersections.
Theorem 6.2 (Generalized Terracini’s Inductive argument).
Consider ${\mathbb{P}}={\mathbb{P}}(1,\ldots,a_{n})$, with coordinate ring $S=k[x_{0},\ldots,x_{n}]$ where $\deg(x_{i})=a_{i}$. Assume $1\leq q\leq r$ and $d\in Z_{+}$ satisfy one of the following inequalities for at least one $i$ in the range $0\leq i\leq n$,
$(n+1)r-s_{d-a_{i}}\leq nq\leq\overline{s}_{d}\qquad\text{ or }\qquad\overline{s}_{d}\leq nq\leq(n+1)r-s_{d-a_{i}},$ (6.1)
where $\overline{S}=S/(x_{i})$ is the coordinate ring of the hyperplane $L=V(x_{i})\cong{\mathbb{P}}(a_{0},\ldots,\hat{a_{i}},\ldots,a_{n})$ and $\overline{s}_{d}=\dim_{k}(\overline{S}_{d})$. If
  1. (1)
    a set of $q$ general double points in $L$ is $\text{AH}_{n-1}(d)$ and
  2. (2)
    the union of a set of $r-q$ general double points in ${\mathbb{P}}$ and a set of $q$ general simple points in $L$ is $\text{AH}_{n}(d-a_{i})$
then a set of $r$ general double points in ${\mathbb{P}}$ is $\text{AH}_{n}(d)$.
Proof.
See Appendix A. ∎
It is notable that Terracini originally came up with Equation 6.1 to study configurations of points in ${\mathbb{P}}^{3}_{k}$. For fixed $d$ and $r$ we refer to the existence of $q$ that satisfies one of the inequalities in Equation 6.1 as “the numerical condition” of Theorem 6.2. In the straight projective space there are infinitely many values of $r$ and $d$ where this condition fails. See [BO08, Section 4]. Even in ${\mathbb{P}}^{2}_{k}$ this numerical condition fails for $d=2,4$ and $r=2,5$, respectively, which are exactly the exceptions in Theorem 1.1.
In Lemma 6.8 we show this numerical condition holds for all $d$ in the case of ${\mathbb{P}}(1,2,3)$ and based on Theorem 7.11 it is the only such space. The main difference with ${\mathbb{P}}^{2}_{k}$ is the different numerics of projective lines in ${\mathbb{P}}(1,2,3)$.
Part (2) of the hypothesis of Theorem 6.2 is satisfied with the aid of Lemma 6.3 which is proved in Appendix A. Originally due to K. Chandler [Cha01, Lemma 3], this lemma is modified to our setting and shows if a certain numerical criterion is satisfied, then one can add simple points all lying on a hyperplane to a set of points that are $\text{AH}_{n}(d)$ while preserving the property. This is made precise in Lemma 6.3.
Lemma 6.3 (Chandler’s Lemma).
Let $I$ be a saturated homogeneous ideal in $S=k[x_{0},\ldots,x_{n}]$ with $\deg(x_{i})=a_{i}$, the homogeneous coordinate ring of ${\mathbb{P}}(a_{0},\ldots,a_{n})$. Assume $l$ is regular on $S/I$. Let $\overline{S}=S/(l)$ and $\deg(l)=i$. Fix $d\in Z_{+}$. Then following are equivalent:
  1. (1)
    A general set $Y_{0}$ of $q$ reduced points in $V(l)$ satisfies
    $H_{d}(S/(I\cap J_{Y_{0}}))=H_{d}(S/I)+q$
  2. (2)
    $H_{d}(S/I)+q\leq H_{d-i}(S/I)+H_{d}(\overline{S})$.
Next we prove $6$ base cases of induction used in proof of Theorem 6.1.
Proposition 6.4.
Let $k$ be an algebraically closed field such that $\operatorname{char}(k)\not\in\{2,3,5\}$. Let $X$ be a set of $r$ double points with $I_{X}$ as its defining ideal. Assume no point of $X$ lies on $V(z)$. Further assume for distinct points $p_{i}$ and $p_{j}$ in $X$ where $p_{i}=[p_{i}^{(1)}:p_{i}^{(2)}:p_{i}^{(3)}]$ and $p_{j}=[p_{j}^{(1)}:p_{j}^{(2)}:p_{j}^{(3)}]$, $p_{i}^{(2)}\neq p_{j}^{(2)}$ for all $i\neq j$. Then $X$ is $\text{AH}_{2}(d)$ for all $0\leq d\leq 5$.
Proof.
By Remark 5.1 it suffices to consider $r=\left\lfloor\frac{1}{3}s_{d}\right\rfloor$ and/or $r=\left\lceil\frac{1}{3}s_{d}\right\rceil$. Note that for $d\leq 5$ we have $s_{d}\leq 5$ and thus $r\leq 2$. Anytime $r=1$, the result follows from Proposition 5.6, whose proof requires the points not to lie on $V(z)$. For $d=4$ or $d=5$ and $r=2$, the result follows from Proposition 5.7, whose proof requires $p_{i}^{(2)}\neq p_{j}^{(2)}$. ∎
We address a subtly in our notion of general and the construction in Theorem 6.2. The condition that no point of $X$ lies on $V(z)$ in Proposition 6.4 does not obstruct the procedure in Theorem 6.2, which requires specializing $q$ points to a hyperplane, $L$. These $q$ points need to be general in $L$; but neither condition $(1)$ or $(2)$ requires $q$ general points in ${\mathbb{P}}(1,2,3)$.
Let $d\geq 6$. Assume any finite set of general double points $X$ is $\text{AH}_{2}(t)$ for $10\leq t\leq d-$ 1. We will use Theorem 6.2, to show for $r=\left\lfloor\frac{1}{3}s_{d}\right\rfloor$ and $r=\left\lceil\frac{1}{3}s_{d}\right\rceil$ that there exists a set of $r$ double points that are $\text{AH}_{2}(d)$. This inductive procedure for $d=14$ is demonstrated in Example 6.7, after which the rest of this section is dedicated to establishing the numerical condition in Lemma 6.8 and condition $(2)$ in Lemma 6.9. To study Equation 6.1 in the context of the ${\mathbb{P}}(1,2,3)$, we introduce the following notation.
Notation 6.5.
Let $S=k[z,u,v]$ denote the coordinate ring of ${\mathbb{P}}(1,2,3)$ where $\deg(z)=1$, $\deg(u)=2$ and $\deg(v)=3$ and recall that $s_{d}=\dim_{k}(S_{d})$. Let $S^{\prime}=k[u,v]$, $S^{\prime\prime}=k[z,v]$ and $S^{\prime\prime\prime}=k[z,u]$ denote the coordinate rings of the projective lines $L_{1}\cong{\mathbb{P}}(2,3)$, $L_{2}\cong{\mathbb{P}}(1,3)$, and $L_{3}\cong{\mathbb{P}}(2,3)$ respectively, where $L_{1}=V(z),L_{2}=V(u),L_{3}=V(v)$ and denote the Hilbert function for each of these rings in degree $d$ by $s^{\prime}_{d},s^{\prime\prime}_{d},s^{\prime\prime\prime}_{d}$, respectively.
These values can be computed explicitly.
Remark 6.6.
For any $d\in{\mathbb{N}}_{0}$ we have
  1. (1)
    $s_{d}=\lfloor d^{2}/12+d/2+1\rfloor$ according to [Sta99, Page 547].
  2. (2)
    $s^{\prime\prime\prime}_{d}=\lfloor d/2\rfloor+1$ cf. Equation 2.7.
  3. (3)
    $s^{\prime\prime}_{d}=\lfloor d/3\rfloor+1$ cf. Equation 2.7.
  4. (4)
    $s^{\prime}_{d}=\begin{cases}\left\lfloor d/6\right\rfloor&d\equiv 1\pmod{6}\\ \left\lfloor d/6\right\rfloor+1&\text{otherwise}.\end{cases}$ cf.  Equation 2.6.
Example 6.7.
We exhibit Terracini’s technique for $d=14$. Since $s_{14}=24$, by Remark 5.1 it suffices to show $8$ double points are $\text{AH}_{2}(14)$. Checking the numerical condition (Equation 6.3 and Equation 6.4) the inequality 3($83(8)-s_{11}=8\leq s^{\prime\prime\prime}_{14}=$ 8 is the only one for which $q$ exists; this corresponds to specializing $4$ double points to ${\mathbb{P}}(1,2)$.
222222222${\mathbb{P}}(1,2)$${\mathbb{P}}(2,3)$${\mathbb{P}}(1,3)$$2$${\mathbb{P}}(1,2)$$1$${\mathbb{P}}(1,2)$$2$
The bottom right part, ${\mathbb{P}}(1,2)$ with $4$ double points, is considered intrinsically in ${\mathbb{P}}(1,2)$ and is $\text{AH}_{1}(14)$ by Proposition 3.2, proving $(1)$ in Theorem 6.2. To justify $(2)$ in Theorem 6.2, we need the bottom left figure, regarded in ${\mathbb{P}}(1,2,3)$, to be $\text{AH}_{2}(11)$. First we show $4$ general double points are $\text{AH}_{2}(11)$ then by Lemma 6.3 their union with 4 simple points on ${\mathbb{P}}(1,2)$ (empty circles) preserves the $\text{AH}_{2}(11)$ property.
Since $s_{11}=16$, to show that any number of points is $\text{AH}_{2}(11)$ it suffices to demonstrate $\left\lfloor\frac{1}{3}s_{11}\right\rfloor=5$ and $\left\lceil\frac{1}{3}s_{11}\right\rceil=6$ points are $\text{AH}_{2}(11)$. We employ Theorem 6.2 again and focus on $5$ points which implies $4$ points are $\text{AH}_{2}(11)$. Checking Equation 6.3 and Equation 6.4, all three inequalities yield a value $q$. We choose $615-s_{8}=5\leq s^{\prime\prime\prime}_{11}=$ 6 which is to specialize $3$ double points to ${\mathbb{P}}(1,2)$.
22222${\mathbb{P}}(1,2)$${\mathbb{P}}(2,3)$${\mathbb{P}}(1,3)$22111${\mathbb{P}}(1,2)$${\mathbb{P}}(1,2)$222
By Proposition 6.4, $2$ general double points are $\text{AH}_{2}(8)$ in ${\mathbb{P}}(1,2,3)$. By Lemma 6.9 their union with $3$ simple points on ${\mathbb{P}}(1,2)$ keeps the configuration $\text{AH}_{2}(8)$. Hence any set of general double points in ${\mathbb{P}}(1,2,3)$ is $\text{AH}_{2}(14)$.
In Lemma 6.8 we establish the numerical condition of Theorem 6.2.
Lemma 6.8.
Let $r=\lfloor\frac{1}{3}s_{d}\rfloor$ or $r=\lceil\frac{1}{3}s_{d}\rceil$ and $d\geq 6$. Then the following equalities can not hold simultaneously.
$\begin{split}&3r-s_{d-1}=s^{\prime}_{d}=\text{positive odd integer},\\ &3r-s_{d-2}=s^{\prime\prime}_{d}=\text{positive odd integer},\\ &3r-s_{d-3}=s^{\prime\prime\prime}_{d}=\text{positive odd integer}\end{split}$ (6.2)
Hence there always exists an even number, say $2q$, such that at least one of the following inequalities holds.
$\begin{split}3r-s_{d-1}&\leq 2q\leq s^{\prime}_{d}\\ 3r-s_{d-2}&\leq 2q\leq s^{\prime\prime}_{d}\\ 3r-s_{d-3}&\leq 2q\leq s^{\prime\prime\prime}_{d}\\ \end{split}$ (6.3)
$\begin{split}s^{\prime}_{d}&\leq 2q\leq 3r-s_{d-1}\\ s^{\prime\prime}_{d}&\leq 2q\leq 3r-s_{d-2}\\ s^{\prime\prime\prime}_{d}&\leq 2q\leq 3r-s_{d-3}\\ \end{split}$ (6.4)
Furthermore, one can pick $q$ such that $1\leq q\leq r$.
Proof.
Assume for the sake of contradiction all equalities in Equation 6.2 hold. The identity $s^{\prime}_{d}+s_{d-1}=s_{d}=3r$ follows from Equation 2.8. This implies $s_{d}$ is divisible by $3$ thus $r=\left\lfloor 1/3s_{d}\right\rfloor=\left\lceil 1/3s_{d}\right\rceil$.
Let $d=6k+i$ where $k\geq 1$ since $d\geq 6$. Plugging $d$ into the formulas from Remark 6.6 gives
  1. I.
    $s_{d}=\lfloor d^{2}/12+d/2+1\rfloor=3k^{2}+(3+i)k+1+\lfloor i^{2}/12+i/2\rfloor$
  2. II.
    $s^{\prime\prime\prime}_{d}=\lfloor d/2\rfloor+1=3k+\lfloor i/2\rfloor+1=$ odd
  3. III.
    $s^{\prime\prime}_{d}=\lfloor d/3\rfloor+1=2k+\lfloor i/3\rfloor+1=$ odd
  4. IV.
    $s^{\prime}_{d}=\begin{cases}k+\left\lfloor i/6\right\rfloor&d\equiv 1\pmod{6}\\ k+\left\lfloor i/6\right\rfloor+1&\text{otherwise}\end{cases}=$ odd
Part III implies $i\neq 3,4,5$, otherwise the $s^{\prime\prime}_{d}$ is not odd. If $i=2$ then IV implies $k$ is even which implies part II, $23k+$ 2, is even. Hence $i\neq 2$. Since $3$ divides $s_{d}$, I implies $i\neq 0$. Finally suppose $i=1$. Part IV implies $k$ must be odd. Hence $13k+$ 1 is even which implies II can not hold. It follows that one of the 6 inequalities in Equation 6.3 or Equation 6.4 holds for any $d\geq 6$. This proves the existence of $q$.
It remains to show one can pick $q$ to satisfy $1\leq q\leq r$. For $1\leq q$, it suffices to show either the lower bound in Equation 6.4 or the upper bound in Equation 6.3, namely $s^{\prime}_{d}$, $s^{\prime\prime}_{d}$ and $s^{\prime\prime\prime}_{d}$ is at least $2$ for $d\geq 6$. For all $i$ and by Equation 2.8
$s^{\prime\prime}_{i}\leq s^{\prime\prime}_{i+1},\;\;s^{\prime\prime\prime}_{i}\leq s^{\prime\prime\prime}_{i+1},$
and $s^{\prime\prime}_{6}=3$ and $s^{\prime\prime\prime}_{6}=4$. Therefore $1\leq q$ if either of the bottom two inequalities in Equation 6.3 or Equation 6.4 holds. If $d=6$ or $d\geq 8$ we also get $s^{\prime}_{d}\geq 2$. For $d=7$, observe that
$13r-s_{6}=\begin{cases}0&\text{ if }r=\left\lfloor\frac{1}{3}s_{7}\right\rfloor\\ 3&\text{ if }r=\left\lceil\frac{1}{3}s_{7}\right\rceil\end{cases}\;,\;s^{\prime}_{7}=$ 1 .
For $r=\left\lfloor\frac{1}{3}s_{7}\right\rfloor$ the inequality $3r-s_{6}\leq s^{\prime}_{7}$ holds true but we do not have a choice of $q$. But we can ignore this and use another inequality because
$43r-s_{d-3}=3\leq s^{\prime\prime\prime}_{d}=$ 4
also holds when $r=\left\lfloor\frac{1}{3}s_{7}\right\rfloor$. Therefore $1\leq q$ for either value of $r$ and $d\geq 6$.
To show $q\leq r$, it is sufficient that either the lower bound in Equation 6.3 or the upper bound in Equation 6.4 is less than or equal to $2r$; that is $3r-s_{d-i}\leq 2r$ for $i\in\{1,2,3\}$. In Equation 6.3 this is sufficient because if $2r$ is less than or equal to the right hand side, we may choose $q$ to be the smallest number such that $2q$ is in the given interval. Note that
$3r-s_{d-i}\leq 2r\Leftrightarrow r\leq s_{d-i}.$
Moreover, considering $r=\left\lceil\frac{1}{3}s_{d}\right\rceil$ and $i=3$ is enough because
$\left\lfloor\frac{1}{3}s_{d}\right\rfloor\leq\left\lceil\frac{1}{3}s_{d}\right\rceil\leq s_{d-3}\leq s_{d-2}\leq s_{d-1}$
Replacing $z=x_{i}$ and $M=S$ in Equation 2.8 for the equivalence ()(*)( ∗ ) yields
$\left\lceil\frac{1}{3}s_{d}\right\rceil\leq s_{d-3}\Leftrightarrow\frac{1}{3}s_{d}<s_{d-3}\underset{*}{\Leftrightarrow}\frac{1}{3}s^{\prime\prime\prime}_{d}+\frac{1}{3}s_{d-3}<s_{d-3}\Leftrightarrow s^{\prime\prime\prime}_{d}<2\cdot s_{d-3}$ (6.5)
Proposition A.2 finishes the proof by showing Equation 6.5 for $d\geq 6$. ∎
Lemma 6.9 proves when Lemma 6.8 is satisfied a set $X$ of $r-q$ double points and $q$ simple points on $L$ is $\text{AH}_{2}(d-i)$ for $i\in\{1,2,3\}$. By Definition 2.17 this means
$H_{d-i}(S/I_{X})=\min\{s_{d-i},3(r-q)+q\}.$
By Lemma 6.8 we know only one of the six inequalities above may be satisfied for any given $d$ and specified values of $r$. Therefore the proof shows any one of the inequalities implies $(2)$ in Theorem 6.2. The proof makes repeated use of Lemma 6.3 and the induction hypothesis.
Lemma 6.9.
Let $d\geq 6$. By Lemma 6.8 at least one of the inequalities
$\begin{split}3r-s_{d-1}&\leq 2q\leq s^{\prime}_{d}\\ 3r-s_{d-2}&\leq 2q\leq s^{\prime\prime}_{d}\\ 3r-s_{d-3}&\leq 2q\leq s^{\prime\prime\prime}_{d}\\ \end{split}$
$\begin{split}s^{\prime}_{d}&\leq 2q\leq 3r-s_{d-1}\\ s^{\prime\prime}_{d}&\leq 2q\leq 3r-s_{d-2}\\ s^{\prime\prime\prime}_{d}&\leq 2q\leq 3r-s_{d-3}\\ \end{split}$ (6.6)
is satisfied and $1\leq q\leq r$.
Let $L$ denote the projective line with defining variable of degree $i$ where an inequality is satisfied. Assume for any $r^{\prime}<r$ a general set of $r^{\prime}$ double points in ${\mathbb{P}}(1,2,3)$ is both $AH_{2}(d-i)$ and $AH_{2}(d-2i)$. Then the union of a set of $r-q$ general double points in ${\mathbb{P}}(1,2,3)$ and a set of $q$ general simple points of $L$ is $\text{AH}_{2}(d-i)$.
Proof.
Denote by $\overline{S}$ the coordinate ring of $L$ where $i\in\{1,2,3\}$ is the corresponding index chosen according to the inequality that holds in Eq. 6.6.
Let $Y$ be a set of $r-q$ general double points in ${\mathbb{P}}(1,2,3)\backslash L$ with defining ideal $I_{q}$ and $J_{q}\subseteq S$ be the defining ideal of $q$ general simple points of $L$ perceived in ${\mathbb{P}}(1,2,3)$. The ideal $I=I_{q}\cap J_{q}$ defines the union of $Y$ with $q$ simple points on $L$.
Case 1: Suppose $3r-s_{d-i}\leq 2q\leq\overline{s}_{d}$. We want to show
$H_{d-i}(S/I)=\min\{s_{d-i},3(r-q)+q\}=3r-2q,$
where the last equality follows from the assumption $3r-s_{d-i}\leq 2q$. Towards this end we apply Lemma 6.3. Condition (2) in this lemma combined with $Y$ being $AH_{2}(d-i)$ and $AH_{2}(d-2i)$ becomes
$\begin{split}H_{d-i}(S/I_{q})+q\leq H_{d-2i}(S/I_{q})+H_{d-i}(\overline{S})\\ \Leftrightarrow 3r-2q\leq\min\{s_{d-2i},3(r-q)\}+\overline{s}_{d-i}.\end{split}$
If $H_{d-2i}(S/I_{q})=s_{d-2i}$, then the right hand side is $s_{d-2i}+\overline{s}_{d-i}=s_{d-i}$ and the inequality $3r-2q\leq s_{d-i}$ is the assumption of this case. On the other hand, if $H_{d-2i}(S/I_{q})=3(r-q)$, we need to show
$3r-2q\leq 3(r-q)+\overline{s}_{d-i}\Leftrightarrow q\leq\overline{s}_{d-i}$
Assumptions of this case imply $q\leq\overline{s}_{d}/2$. Lemma A.3 illustrates
$\overline{s}_{d}/2\leq\overline{s}_{d-i},$ (6.7)
which establishes the desired inequality $q\leq\overline{s}_{d-i}$.
Since condition (2) of Lemma 6.3 holds, by the equivalent condition (1) of this result the ideal $I=I_{q}\cap J_{q}$ satisfies
$H_{d-i}(S/I)=H_{d-i}(S/I_{q})+q.$
By hypothesis $Y$ is $\text{AH}_{2}(d-i)$ and further $3(r-q)<3r-2q\leq s_{d-i}$. Hence the displayed identity leads to the desired conclusion
$H_{d-i}(S/I_{q})=3(r-q)=\min\{s_{d-i},3(r-q)\}.$
Case 2: Suppose $\overline{s}_{d}\leq 2q\leq 3r-s_{d-i}$. We wish to demonstrate
$H_{d-i}(S/I)=\min\{s_{d-i},3r-2q\}=s_{d-i}.$
Since $Y$ is $AH_{2}(d-i)$ we have
$H_{d-i}(S/I_{q})=\min\{s_{d-i},3(r-q)\}.$
If $H_{d-i}(S/I_{q})=s_{d-i}$, we obtain
$[I_{q}]_{d-i}=0\Rightarrow[I]_{d-i}=[I_{q}]_{d-i}\cap[J_{q}]_{d-i}=0\Rightarrow H_{d-i}(S/I)=s_{d-i}.$
Suppose $H_{d-i}(S/I_{q})=3(r-q)<s_{d-i}$. Set $m=s_{d-i}-3(r-q)$ and note $m\leq q$ by assumption of this case. Let $J_{m}$ be the defining ideal of $m$ among the chosen simple points on $L$, $J_{q-m}$ the ideal defining the remaining simple points, and $F_{m}=I_{q}\cap J_{m}$. It is enough to show $(F_{m})_{d-i}=0$. Then
$[I]_{d-i}=[I_{q}]_{d-i}\cap[J_{m}]_{d-i}\cap[J_{q-m}]_{d-i}=[F_{m}]_{d-i}\cap[J_{q-m}]_{d-i}=0$
implies $H_{d-i}(S/I)=s_{d-i}$. Equivalently we want to show
$H_{d-i}(S/F_{m})=H_{d-i}(S/I_{q})+m=s_{d-i}.$ (6.8)
By Lemma 6.3, Equation 6.8 is true if and only if
$s_{d-i}\leq H_{d-2i}(S/I_{q})+\overline{s}_{d-i}=\min\{s_{d-2i},3(r-q)\}+\overline{s}_{d-i},$ (6.9)
where the last equality follows from $Y$ being $AH_{2}(d-2i)$. It suffices to show $q\leq\overline{s}_{d-i}$. If true, we use the assumption $s_{d-i}\leq 3r-2q$ to conclude $s_{d-2i}=s_{d-i}-\overline{s}_{d-i}\leq 3(r-q)$. Then the right hand side of Equation 6.9 becomes $s_{d-i}$ and the inequality claimed therein holds as an equality. Therefore we have established Equation 6.8, which settles the claim as explained above. To verify the remaining inequality $q\leq\overline{s}_{d-i}$, observe that
$q\leq r\leq\left\lceil\frac{1}{3}s_{d}\right\rceil\text{ and }s_{d-3}\leq s_{d-2}\leq s_{d-1}.$
The inequality $\left\lceil\frac{1}{3}s_{d}\right\rceil\leq s_{d-3}$ follows from Eq. 6.5 and Proposition A.2. ∎
7. Exceptions
7.1. An interpolation bound for ${\mathbb{P}}(1,b,c)$
Let $k$ be an algebraically closed field of characteristic 00. The main goal of this section to give an interpolation bound for weighted projective planes of the form ${\mathbb{P}}(1,b,c)$ in Theorem 7.7. That is to find $t\in{\mathbb{N}}$ such that for all $d\geq t$ we can guarantee that any set of $r$ general double points $X$ is $\text{AH}_{2}(d)$, or not exceptional, i.e.
$H_{d}(S/I_{X})=\min\{s_{d},3r\}.$
In general double point interpolation in ${\mathbb{P}}(1,b,c)$ is drastically different than in ${\mathbb{P}}(1,2,3)$ and ${\mathbb{P}}_{k}^{2}$ as indicated by Example 7.2.
Definition 7.1.
For a set of $r$ general double points $X\subseteq{\mathbb{P}}(a,b,c)$ we define deficiency of $X$ in degree $d$ as
$D(X,d):=\min\{s_{d},3r\}-H_{d}(S/I_{X})$
By Remark 5.2 when at least one weight is $1$ we have $D(X,d)\geq 0$. Further using this terminology, Theorem 1.1 states that $D(X,d)=0$ in ${\mathbb{P}}_{k}^{2}$ with exceptions when $d=2,\;r=2$ and $d=4,\;r=5$ and deficiency is exactly $1$ in each case. Computations with Macaulay2 [GS] illustrate even in ${\mathbb{P}}(1,b,c)$ this is not the case.
Example 7.2.
We demonstrate some examples of exceptional cases in ${\mathbb{P}}(1,b,c)$.
Weighted ${\mathbb{P}}^{2}$#\## of general points Deficiency
${\mathbb{P}}(1,5,9)$ 3 $D(X,d)=1$ for $20\leq d\leq 22$
${\mathbb{P}}(1,5,26)$ 2 $D(X,d)=2$ for $d=25$ $D(X,d)=1$ for $d=20-24,26-30$
${\mathbb{P}}(1,4,57)$ 4 $D(X,d)=1$ for $d=32-35,65-68$ $D(X,d)=2$ for $d=36-39,61-64$ $D(X,d)=3$ for $d=40-43,57-60$ $D(X,d)=4$ for $d=44-56$
Table 1. Exceptional Cases in ${\mathbb{P}}(1,b,c)$
Example 7.2 shows that even when the ideal of points in a weighted projective plane are defined by complete intersections, their behavior can be different than points in the usual ${\mathbb{P}}^{2}_{k}$. Our Macaulay2 experimentation suggests by increasing the difference between the two largest weights one can arbitrarily increase $D(X,d)$ for a suitable number of general double points.
Problem 7.3.
If $d\in{\mathbb{N}}$ is allowed to vary, is $D(X,d)$ bounded in any weighted projective plane?
Proof of Theorem 7.7 mimics Terracini’s proof of Theorem 1.1 for $n=2$; [Ter11]. It relies on two two prelimenary results, Theorem 7.5 and Lemma 7.6. We introduce the necessary terminology to state these results.
Definition 7.4.
Let $S=k[x_{0},\ldots,x_{n}]$ be the coordinate ring of ${\mathbb{P}}(a_{0},\ldots,a_{n})$, $N=s_{d}-1$, $m_{0},\ldots,m_{N}$ enumerate the monomials in $S_{d}$ and $I$ be the kernel of the $k$-algebra homomorphism
$\begin{split}\theta:k[y_{0},\ldots,y_{N}]&\to k[x_{0},\ldots,x_{n}]\\ y_{i}&\to m_{i}.\end{split}$
The projective variety
$V^{d}=\operatorname{Proj}(k[y_{0},\ldots,y_{N}]/I)\subset{\mathbb{P}}^{N}_{k}$
is called the $d$-th Veronese variety of ${\mathbb{P}}(a_{0},\ldots,a_{n})$. Geometrically $\theta$ induces a map of projective varieties
$\psi:{\mathbb{P}}(a_{0},\ldots,a_{n})\dashrightarrow V^{d},\,\psi(p)=[m_{0}(p):\cdots:m_{N}(p)]$
defined on $U=\{p\in\operatorname{Proj}(S)\;|\;\theta((y_{0},\ldots,y_{N}))\not\subseteq p\}$.
Note that unlike the straight projective space, the Veronese map is only partially defined on ${\mathbb{P}}(a_{0},\ldots,a_{n})$. Henceforth the notion of a set of general points $X$ includes $X\subseteq U$.
By $\sigma_{r}(V^{d})$ we denote the $r$-th secant variety to the image of the Veronese map. For a point $p\in V^{d}$, we let $T_{p}V^{d}$ denote the tangent space of $V^{d}$ at p. For a precise definition of each see Definition B.7 and Definition B.3.
Theorem 7.5.
For a fixed $d\geq a_{n}$ and a set of $r$ general double points $X\subseteq{\mathbb{P}}(1,a_{1},\ldots,a_{n})$, the following are equivalent:
  1. (1)
    ${H_{d}(S/I_{X})=\min\{s_{d},r(n+1)\}}$
  2. (2)
    $\dim_{k}(\sigma_{r}(V^{d}))=\min\{s_{d}-1,r(n+1)-1\}$.
Proof.
See Appendix B. ∎
Lemma 7.6.
(Second Terracini lemma) Let $U$ be as in Definition 7.4. Let $X=\{2p_{i}\}_{i=1}^{r}\subseteq U\subseteq{\mathbb{P}}(a_{0},\ldots,a_{n})$ be a set of double points with defining ideal $I_{X}$ and fix $d\geq a_{n}$. We identify the reduced points $p_{i}$ with their images on $V^{d}$ according to the Veronese map.
Assume that $H_{d}(S/I_{X})<\min\{s_{d},r(n+1)\}$. Then there is a positive dimensional variety $C\subseteq V^{d}$ through $p_{1},\ldots p_{r}$ such that if $p\in C$ then $T_{p}V^{d}\subseteq<T_{p_{1}}V^{d},\ldots,T_{p_{r}}V^{d}>$. In particular, by Proposition B.5, every hypersurface of degree $d$ which is singular at $p_{i}$ is also singular along $C$.
Proof.
See [BO08, Lemma 2.3]
We note the proof of Lemma 7.6 in [BO08] is completely independent of weights and holds in the setting of weighted projective space.
Theorem 7.7.
A set of $r$ general double points $X\subseteq{\mathbb{P}}(1,b,c)$ is $\text{AH}_{2}(d)$,
$H_{d}(S/I_{X})=\min\{s_{d},3r\},$ (7.1)
for all $d\geq 10c$. Additionally, Equation 7.1 holds if $d\geq 6c$ and $\left\lfloor\frac{2c}{b}\right\rfloor\geq 5$.
Proof.
Let $X$ be a set of $r$ general double that does not satisfy Equation 7.1 for some degree $d\geq c$. There exists $F\in(I_{X})_{d}$ where $F\neq 0$. By Lemma 7.6, there is a positive dimensional variety $C\subseteq V^{d}$ through the image of the points in $V^{d}$. By abuse of notation we now identify $C$ with its preimage under the Veronese map. By the last part of Lemma 7.6, $F$ is singular along $C$. Hence $F$ contains a double component through $X$, namely $F=C^{2}H$. Let $\deg(C)=l$. Since one of the weights is $1$, Remark 5.2 (3) implies
$2l\leq d\Rightarrow l\leq\left\lfloor\frac{d}{2}\right\rfloor\Rightarrow s_{l}\leq s_{\left\lfloor\frac{d}{2}\right\rfloor}.$ (7.2)
By Lemma 7.6, $C$ vanishes on the reduced set $\{p_{i}\}_{i=1}^{r}$. From Theorem 4.1 and Remark 5.1 we get
$\left\lfloor\frac{1}{3}s_{d}\right\rfloor\leq r\leq s_{l}-1.$ (7.3)
Putting Equation 7.2 and Equation 7.3 together we acquire
$\left\lfloor\frac{1}{3}s_{d}\right\rfloor\leq s_{\left\lfloor\frac{d}{2}\right\rfloor}-1.$ (7.4)
Hence for values of $d$ where Equation 7.4 does not hold, any set $X$ satisfies Equation 7.1. In Lemma 7.8 we show that the inequalities involving $d$ given in the hypothesis force Equation 7.4 to fail. This completes the proof. ∎
Lemma 7.8.
If either of the following is true then $\left\lfloor\frac{1}{3}s_{d}\right\rfloor\geq s_{\left\lfloor\frac{d}{2}\right\rfloor}$ holds.
  1. (1)
    $d\geq 10c$
  2. (2)
    $d\geq 6c$ and $\left\lfloor\frac{2c}{b}\right\rfloor\geq 5$
Proof.
We employ an interpretation of the Hilbert function of a nonstandard graded polynomial ring as lattice point counting.
Observe $s_{d}$ is the number of lattice points on and inside the triangle $T$ with vertices $(0,0),\left(0,\frac{d}{c}\right),\left(\frac{d}{b},0\right)$. Recall $s_{d}$ is the number solutions to $r_{0}+br_{1}+cr_{2}=d$ where $r_{i}\in{\mathbb{N}}_{0}$. Which are in bijection with lattice points $(r_{1},r_{2})$ where $br_{1}+cr_{2}\leq d$, $r_{1}\geq 0$ and $r_{2}\geq 0$. The latter is the number of lattice points on and inside of $T$ which we now divide into smaller triangles.
[…]
Triangle $T_{1}$ has vertices at $(0,0),\left(0,\frac{\left\lfloor\frac{d}{2}\right\rfloor}{c}\right),\left(\frac{\left\lfloor\frac{d}{2}\right\rfloor}{b},0\right)$. Our goal is to impose conditions on $b$ and $c$ so that $T_{1}$, $T_{2}$ and $T_{3}$ have small intersections and $T_{4}$ contains enough lattice points to imply the desired inequality. By #$\#T$ we refer to the number of lattice points on and inside the triangle $T$. Note that
  • #$\#T_{1}=s_{\left\lfloor\frac{d}{2}\right\rfloor}$,
  • $T_{2}$ and $T_{3}$ are translates of $T_{1}$ by integer vectors hence
    #$\#T_{2}=\#T_{3}=\#T_{1}=s_{\left\lfloor\frac{d}{2}\right\rfloor}.$
We make the following observations
  1. (1)
    $T_{1}\cap T_{2}=\left\{\left(\left\lfloor\frac{d}{2b}\right\rfloor,0\right)\right\}$.
  2. (2)
    $T_{2}\cap T_{3}=\left\{\left(\left\lfloor\frac{d}{2b}\right\rfloor,\left\lfloor\frac{d}{2c}\right\rfloor\right)\right\}$
  3. (3)
[…]
  4. (4)
    $T_{1}\cap T_{2}\cap T_{3}=\emptyset$.
Item $(4)$ follows from $(2)$ once we check the point $\left(\left\lfloor\frac{d}{2b}\right\rfloor,\left\lfloor\frac{d}{2c}\right\rfloor\right)$ is not in $T_{1}$. The slope of the hypotenuse of $T_{1}$ is $\frac{-b}{c}$, so the point is in $T_{1}$ if and only if $\left\lfloor\frac{d}{2c}\right\rfloor\leq\frac{b}{c}<1$. But this can’t happen since $d\geq 2c$ implies $\left\lfloor\frac{d}{2c}\right\rfloor\geq 1$.
Using $(1)-(4)$ above and denoting by #$\#\accentset{\circ}{T}_{4}$ the number of lattice points in the interior of $T_{4}$, we obtain
$s_{d}$ =#$=\#T$
#$\geq\#T_{1}+\#T_{2}+\#T_{3}-\#(T_{1}\cap T_{2})-\#(T_{2}\cap T_{3})-\#(T_{1}\cap T_{3})+\#\accentset{\circ}{T}_{4}$
$\geq 3s_{\left\lfloor\frac{d}{2}\right\rfloor}-3-\left\lfloor\frac{c}{b}\right\rfloor+\#\accentset{\circ}{T}_{4}.$ (7.5)
We now estimate #$\#\accentset{\circ}{T}_{4}$. Let $(p,q)$ be the point of intersection between the hypotenuse of $T_{1}$ and the line $x=\left\lfloor\frac{d}{2b}\right\rfloor$. Using the figure above one can show
$q<i<\left\lfloor\frac{d}{2c}\right\rfloor\text{ and }\frac{\left\lfloor\frac{d}{2}\right\rfloor}{b}-\frac{ic}{b}<t<\left\lfloor\frac{d}{2b}\right\rfloor\iff(t,i)\in\accentset{\circ}{T}_{4}$ (7.6)
For a fixed $i$, we obtain $\left\lfloor\frac{ic}{b}\right\rfloor-1\geq i-1$ lattice points in #$\#\accentset{\circ}{T}_{4}$. Let $d\geq 10c$. From $\left\lfloor\frac{d}{2c}\right\rfloor\geq 5$ and Equation 7.6 we get $i\leq 4$. Since $t_{5}$ is similar to $T$ and $\frac{\left\lfloor\frac{d}{2}\right\rfloor}{b}-\left\lfloor\frac{d}{2b}\right\rfloor\leq 1$ we get $q<1$, hence $1\leq i$. Plugging $1\leq i\leq 4$ in Equation 7.6 one gets
#$\#\accentset{\circ}{T}_{4}\geq\left\lfloor\frac{c}{b}\right\rfloor+5.$
Plugging this into Section 7.1 yields the following and proves $(1)$.
$s_{d}\geq 3s_{\left\lfloor\frac{d}{2}\right\rfloor}-\left\lfloor\frac{c}{b}\right\rfloor-3+\left\lfloor\frac{c}{b}\right\rfloor+5=3s_{\left\lfloor\frac{d}{2}\right\rfloor}+2.$
Suppose $d\geq 6c$ and $\left\lfloor\frac{2c}{b}\right\rfloor\geq 5$. By $\left\lfloor\frac{d}{2c}\right\rfloor\geq 3$ and Equation 7.6 we get
#$\#\accentset{\circ}{T}_{4}\geq\left\lfloor\frac{c}{b}\right\rfloor-1+\left\lfloor\frac{2c}{b}\right\rfloor-1.$
By $\left\lfloor\frac{2c}{b}\right\rfloor\geq 5$ and Section 7.1 we arrive at the desired conclusion
$s_{d}\geq 3s_{\left\lfloor\frac{d}{2}\right\rfloor}-\left\lfloor\frac{c}{b}\right\rfloor-3+\left\lfloor\frac{c}{b}\right\rfloor-1+\left\lfloor\frac{2c}{b}\right\rfloor-1\geq 3s_{\left\lfloor\frac{d}{2}\right\rfloor}.\qed$
7.2. Exceptions in ${\mathbb{P}}(1,b,c)$.
So far we have shown exceptional sets of points in ${\mathbb{P}}(1,b,c)$ only happen when $d<10c$. As a porism of Theorem 7.7 we get Corollary 7.10 which fully classifies exceptional cases of ${\mathbb{P}}(1,b,c)$ in degrees $d$ where $s_{d}$ is divisible by $3$. Proposition 7.9 gives a sufficient condition for a set of points to fail Equation 7.1.
Proposition 7.9.
Fix $d\in{\mathbb{N}}$. Let $X\subseteq{\mathbb{P}}(1,a_{1},\ldots,a_{n})$ be a set of $r$ general simple points. Let $J_{X}$ and $I_{X}$ be the defining ideal of $X$ and $2X$ respectively; with $2X$ denoting the set of double points supported on $X$. If $r<s_{\left\lfloor\frac{d}{2}\right\rfloor}$ and $(n+1)r\geq s_{d}$ then $X$ is not $\text{AH}_{n}(d)$.
Proof.
Let $e=\left\lfloor\frac{d}{2}\right\rfloor$. By Theorem 4.1 and the assumption $r<s_{e}$, there exists $F\in(J_{X})_{e}$ where $F\neq 0$. The proof follows from the following simple observation.
$F^{2}\in(J^{2}_{X})_{2e}\subseteq(I_{X})_{2e}$
However the expected dimension of $(I_{X})_{2e}=0$ because
$(n+1)\cdot r\geq s_{d}\underset{*}{\geq}s_{2e},$
[…]
Corollary 7.10.
Let $d\in{\mathbb{N}}$ be such that $s_{d}=3r$. Then a set $X\subseteq{\mathbb{P}}(1,b,c)$ of $r$ general double points is not $\text{AH}_{2}(d)$ if and only if $r<s_{\left\lfloor\frac{d}{2}\right\rfloor}$.
Proof.
Let $X$ be as defined in the hypothesis. Using Proposition 7.9 for ()($\Rightarrow$)( ⇒ ) and Equation 7.4 for ()($\Leftarrow$)( ⇐ ), which is a consequence of not being $\text{AH}_{2}(d)$ as shown in the proof of Theorem 7.7, we get
$\begin{split}&r=\frac{1}{3}s_{d}\leq s_{\left\lfloor\frac{d}{2}\right\rfloor}-1<s_{\left\lfloor\frac{d}{2}\right\rfloor}\\ &\iff\text{ X does does not impose independent conditions in degree $d$.}\qed\end{split}$
Using Corollary 7.10 one can prove ${\mathbb{P}}(1,2,3)$ is the only space of the form ${\mathbb{P}}(1,b,c)$ with no exceptional sets of general double points.
Theorem 7.11.
Let $X\subseteq{\mathbb{P}}(1,b,c)$ be a finite set of general double points. Then $X$ is $\text{AH}_{2}(d)$ for all $d\in{\mathbb{N}}$ if and only if $b=2$ and $c=3$.
Proof.
()($\Rightarrow$)( ⇒ ) The proof is broken into the cases $b=1$, $b=2$ and $b\geq 3$ where $r$ denotes the number of points in $X$. If $b=1$, $c=2$ and $r=3$ then $X$ is exceptional because $s_{2}=4$ and by Theorem 4.1 there exists a quadratic form $F$ through the $3$ points resulting in $F^{2}\in[I_{X}]_{4}$. But the fact that $s_{4}=9$ yields
$H_{S/I_{X}}(4)\leq 8<9=\min\{9,9\}.$
If $c\geq 3$ by Proposition 5.4 one general double point is exceptional.
For $b=2$ and $r=1$, Proposition 5.6 implies $c<4$ and further $2<c$ by Remark 2.4, therefore $c=3$ and we are done. Let $b\geq 3$. We contradict the hypothesis by proving it can not hold for $r=1,2,3$ points and all $d$ simultaneously. When $r=2$, Proposition 5.7, which uses the assumption $b\neq 2$, implies $\frac{3}{2}b<c<2b$ or $2b<c<3b$. By Proposition 5.6 the latter would contradict $1$ point being $\text{AH}_{2}(d)$ for all $d$ therefore $\frac{3}{2}b<c<2b$. This criteria on $c$ results in $s_{4b}=9$ and $s_{2b}=4$ thus Corollary 7.10 implies $3$ points are not $\text{AH}_{2}(d)$. Theorem 6.1 yields the backwards direction. ∎
8. Open problems
We end this paper with a list of open problems, some of which have been stated through out the paper.
  1. (1)
    With possible restrictions on the weights, find explicit generators for the defining ideal of a general point in ${\mathbb{P}}(a_{0},\ldots,a_{n})$ for $n\geq 3$.
  2. (2)
    Using the notation introduced in Section 7, is $D(X,d)$ bounded uniformly (independently of $d$) in any weighted projective plane?
  3. (3)
    Let ${\mathbb{P}}(1,a_{1},\ldots,a_{n})$ and $g=\operatorname{lcm}(a_{1},\ldots,a_{n})$. One can show if $g\mid d$ then the Veronese map in Definition 7.4 becomes an embedding. Compute the minimal free resolution of the ideal defining the Veronese variety $V^{d}$ for such $d$.
  4. (4)
    The well-known SHGH conjecture, see [CHHVT20], states the exceptional degrees for a set $X\subseteq{\mathbb{P}}^{2}$ where $X$ contains points with mixed multiplicity. State the analogue of this conjecture for ${\mathbb{P}}(1,b,c)$.
  5. (5)
    We suspect that méthode d’Horace is applicable in ${\mathbb{P}}(1,a_{1},\ldots,a_{n})$ for $d\geq\operatorname{lcm}(a_{1},\ldots,a_{n})$. But we could not adapt the details of the modern proof due to Chandler, in particular the use curvilinear subschemes, in the setting of weighted projective space. We leave it as a conjecture that méthode d’Horace, or a simple modification of it, applies in weighted projective space.
Appendix A Proofs related to Section 6
This section contains proofs for several foundational results used in Section 6 adapted from [HM21] and [Cha01] as well as proofs for some numerical inequalities we employ in that section. We use 2.12. Our proof of Theorem 6.2 mimics the one presented in [HM21]; with expansion on two details that were not clear to us. The proof of Lemma 6.9 gives an accessible proof of [Cha01, Lemma 3] when all the weights are $1$.
Let $\mathfrak{m}$ be the homogeneous maximal ideal of a graded ring and let $I$ be any ideal. The ideal $(I:\mathfrak{m}^{\infty})$ is called the saturation of $I$, denoted $I^{\text{sat}}$.
Lemma A.1 (Generalized Castelnuovo’s Inequality).
Let $S$ be a not necessarily standard graded polynomial ring, $I$ a homogeneous ideal and $l$ a form of degree $k$ in $S$. Set $\tilde{I}=I:l$, $\overline{S}=S/(l)$ and $\overline{I}=(I+(l))/(l)$, the ideal of $I$ in $\overline{S}$. Then
$H_{d}(S/I)\geq H_{d-k}(S/\tilde{I})+H_{d}(\overline{S}/(\overline{I})^{\text{sat}}),$
with equality if and only if $\overline{I}$ is saturated in $\overline{R}$.
Proof.
See [HM21, Lemma 2.2]. ∎
We note the assumption on the ambient space in Theorem 6.2 allows us to conclude the defining ideal of a simple point is a complete intersection (Corollary 2.9, Proposition 2.10) which is implicitly used through out the proof. For a set of reduced points $X$ with defining ideal $J_{X}$, let $2X$ denote the set of double points supported on $X$ where $I_{X}=J^{(2)}_{X}$; see Definition 2.15.
Theorem 6.2 (Generalized Terracini’s Inductive argument).
Consider ${\mathbb{P}}={\mathbb{P}}(1,a_{1},\ldots,a_{n})$, with the coordinate ring $S=k[x_{0},\ldots,x_{n}]$ where $\deg(x_{i})=a_{i}$. Assume $1\leq q\leq r$ and $d\in{\mathbb{N}}$ satisfy one of the following inequalities for at least one $i$ in the range $0\leq i\leq n$,
$(n+1)r-s_{d-a_{i}}\leq nq\leq\overline{s}_{d}\qquad\text{ or }\qquad\overline{s}_{d}\leq nq\leq(n+1)r-s_{d-a_{i}},$ (A.1)
where $\overline{S}=S/(x_{i})$ is the coordinate ring of the hyperplane $L=V(x_{i})\cong{\mathbb{P}}(a_{0},\ldots,\hat{a_{i}},\ldots,a_{n})$ and $\overline{s}_{d}=\dim_{k}(\overline{S}_{d})$. If
  1. (1)
    a set of $q$ general double points in $L$ is $\text{AH}_{n-1}(d)$ and
  2. (2)
    the union of a set of $r-q$ general double points in ${\mathbb{P}}$ and a set of $q$ general simple points in $L$ is $\text{AH}_{n}(d-a_{i})$
then a set of $r$ general double points in ${\mathbb{P}}$ is $\text{AH}_{n}(d)$.
Proof.
[…]
For any $d$, the definition of Hilbert function and Definition 2.18 imply $H_{d}(S/I)\leq s_{d}$ and $H_{d}(S/I)\leq r(n+1)$ hence it suffices to show $H_{d}(S/I)\geq\min\{s_{d},r(n+1)\}$. Since $Y_{1}\subseteq L$ and $\𝐿Y_{2}\subseteq{\mathbb{P}}\backslash L$ we get
$\tilde{I}:=(I:x_{i})=(J^{(2)}_{Y_{1}}\cap J^{(2)}_{Y_{2}}:x_{i})=(J^{(2)}_{Y_{1}}:x_{i})\cap(J^{(2)}_{Y_{2}}:x_{i})=J_{Y_{1}}\cap J^{(2)}_{Y_{2}}.$
To justify the final equality we will show $(J^{(2)}_{Y_{1}}:x_{i})\subseteq J_{Y_{1}}$. Observe
$s\in(J^{(2)}_{Y_{1}}:x_{i})=(\displaystyle\bigcap_{j=1}^{q}J^{(2)}_{p_{j}}:x_{i})=\bigcap_{j=k}^{q}(J^{2}_{p_{j}}:x_{i})\Rightarrow s\cdot x_{i}\in\bigcap_{i=k}^{q}J^{2}_{p_{j}}.$
We want to show $s\in J_{p_{j}}$ for all $j$.
Fix a $j$ and note $s\cdot x_{i}=0$ in $\operatorname{gr}_{J_{p_{j}}}S$, where $s\in S/J_{p_{j}}$ and $x_{i}\in J_{p_{j}}/J^{2}_{p_{j}}$, but clearly $x_{i}\not\in J^{2}_{p_{j}}$. Using the generators of $J_{p_{j}}$ from Corollary 2.9 or Proposition 2.10, we can show $S_{J_{p_{j}}}$ is a regular ring, hence $\operatorname{gr}_{J_{P_{j}}}S$ is a domain. Therefore $s\in J_{p_{j}}$ for all $j$. Note $\tilde{I}$ defines a set of $q$ general simple points in $L$ and $r-q$ general double points in ${\mathbb{P}}$.
Claim: $(\overline{I})^{\text{sat}}={\overline{J_{Y_{1}}}}^{(2)}$. Observe that $\overline{I}\subseteq\overline{J_{Y_{1}}}$. Since $\overline{S}/\overline{J_{Y_{1}}}$ is a one-dimensional ring, $(\overline{I})^{(\text{sat})}$ is the intersection of minimal components of $\overline{I}$. These are minimal primes containing $I=J_{Y_{1}}^{(2)}\cap J_{Y_{2}}^{(2)}$ and $x_{i}$. Since no minimal prime of $J_{Y_{2}}$ contains $x_{i}$, we get $\operatorname{Min}\{(I,x_{i})\}=\operatorname{Min}\{J_{Y_{1}}\}$. For any ${\mathfrak{p}}\in\operatorname{Min}\{(I,x_{i})\}$ we get
$\begin{split}(I,x_{i})^{\text{sat}}_{\mathfrak{p}}=(I,x_{i})_{\mathfrak{p}}=(J_{Y_{1}}^{(2)}\cap J_{Y_{2}}^{(2)},x_{i})_{\mathfrak{p}}&=(J_{Y_{1}}^{(2)},x_{i})_{\mathfrak{p}}\cap(J_{Y_{2}}^{(2)},x_{i})_{\mathfrak{p}}\\ &=(J_{Y_{1}}^{(2)},x_{i})_{\mathfrak{p}}=(J_{Y_{1}}^{(2)},x_{i})^{\text{sat}}_{\mathfrak{p}}.\end{split}$
This proves the claim. By Lemma A.1 we get
$H_{d}(S/I)\geq H_{d-a_{i}}(S/\tilde{I})+H_{d}(\overline{S}/\overline{I}^{\text{sat}})=\min\{s_{d-a_{i}},(n+1)(r-q)+q\}+\min\{\overline{s}_{d},nq\}$ (A.2)
where the last equality follows from assumptions $(1)$ and $(2)$.
If $(n+1)r-s_{d-a_{i}}\leq nq\leq\overline{s}_{d}$ holds, the right hand side of Equation A.2 becomes $(n+1)r$. If $\overline{s}_{d}\leq nq\leq(n+1)r-s_{d-a_{i}}$ holds, the right hand side of Equation A.2 is $s_{d-a_{i}}+\overline{s}_{d}=s_{d}$ by Remark 2.14. Note if $H_{d}(S/I)$ is equal to $(n+1)r$ or $s_{d}$ then that value is the minimum of the two. ∎
Lemma 6.3 (Chandler’s Lemma).
Let $I$ be a saturated homogeneous ideal in $S=k[x_{0},\ldots,x_{n}]$ with $\deg(x_{i})=a_{i}$, the homogeneous coordinate ring of ${\mathbb{P}}(a_{0},\ldots,a_{n})$. Assume $l$ is regular on $S/I$. Let $\overline{S}=S/(l)$ and $\deg(l)=i$. Fix $d\in Z_{+}$. Then following are equivalent:
  1. (1)
    A general set $Y_{0}$ of $q$ reduced points in $V(l)$ satisfies
    $H_{d}(S/(I\cap J_{Y_{0}}))=H_{d}(S/I)+q.$ (A.3)
  2. (2)
    $H_{d}(S/I)+q\leq H_{d-i}(S/I)+H_{d}(\overline{S})$.
Proof.
$(1)\Rightarrow(2)$ Since $1\cdot l\in J_{Y_{0}}$ and $l$ is regular on $S/I$ we get
$(I\cap J_{Y_{0}}:(l))=(I:(l))\cap(J_{Y_{0}}:(l))=(I:(l))\cap S=(I:(l))=I$
Using this equality and the restriction sequence
$0\to S/(I\cap J_{Y_{0}}:l)(-i)\xrightarrow{\cdot l}S/(I\cap J_{Y_{0}})\twoheadrightarrow S/(I\cap J_{Y_{0}},l)\to 0$
we get
$H_{d}(S/I\cap J_{Y_{0}})=H_{d-i}(S/I)+H_{d}(S/(I\cap J_{Y_{0}},l))\leq H_{d-i}(S/I)+H_{d}(\overline{S}).$
$(2)\Rightarrow(1)$ Suppose there exists a positive integer $q$ satisfying the inequality in (2). We proceed by induction on $j$ where
$0\leq j\leq H_{d-i}(S/(I:l))+H_{d}(\overline{S})-H_{d}(S/I)$
to construct a set $Y(j)$ of $j$ reduced points in $V(l)$ that satisfies
$H_{d}(S/(I\cap J_{Y(j)}))=H_{d}(S/I)+j.$ (A.4)
If $j=0$, let $Y(0)=\emptyset$. For $0\leq j<q$ suppose there is a set $Y(j)\subseteq V(l)$ that satisfies Equation A.4. Since $j+1\leq q$ assumption (2) gives
$j+1\leq q\leq H_{d-i}(S/(I:l))+H_{d}(\overline{S})-H_{d}(S/I)$
from which it follows
$H_{d}(S/(I\cap J_{Y(j)}))-H_{d-i}(S/(I:l))\leq H_{d}(\overline{S})-1.$
Recall that $(I\cap J_{Y(j)}:l)=(I:l)$ hence
$H_{d}(S/(I\cap J_{Y(j)}))-H_{d-i}(S/(I\cap J_{Y(j)}:l))\leq H_{d}(\overline{S})-1$ (A.5)
Using Equation A.5 and the restriction sequence
$0\to S/(I\cap J_{Y(j)}:(l))(-i)\xrightarrow{\cdot l}S/(I\cap J_{Y(j)})\twoheadrightarrow S/(I\cap J_{Y(j)}+l)\to 0$
We get
$1\leq H_{d}(\overline{S})-H_{d}(S/(I\cap J_{Y(j)}+l))=H_{d}((I\cap J_{Y(j)}+l)/(l))$
Therefore there exists $F\in[I\cap J_{Y(j)}+l]_{d}\setminus[(l)]_{d}$. Let $p\in V(l)\setminus V(F)$. Such $p$ exists because $F\not\in(l)$. Further note
$F\in((I\cap J_{Y(j)}+l)/l)\Rightarrow V(F)\supseteq V(I\cap J_{Y(j)})\cap V(l)\supseteq Y(j).$
Since $p\not\in V(F)$ we get $Y(j+1)=Y(j)\cup\{p\}$ has $j+1$ distinct points. We obtain
$H_{d}(S/(I\cap J_{Y(j+1)}))\geq H_{d}(S/(I\cap J_{Y(j)}))+1=H_{d}(S/I)+j+1,$
where the first inequality follows since $F\in[I\cap J_{Y(j)}]_{d}\setminus[I\cap J_{Y(j+1)}]_{d}$ and the last equality follows from the induction hypothesis. Finally, the exact sequence below yields the other inequality.
$0\to\cfrac{S}{I\cap J_{Y(j+1)}}\to S/I\oplus S/J_{Y(j+1)}\to S/I+J_{Y(j+1)}\to 0$
$H_{d}(S/(I\cap J_{Y(j+1)}))\leq H_{d}(S/I)+H_{d}(S/J_{Y(j+1)})\leq H_{d}(S/I)+j+1.$ (A.6)
We have shown the set $Y(q)$ satisfies Equation A.4. By semicontinuity of Hilbert functions a general set $Y_{0}$ of $q$ points then satisfies $H_{d}(S/(I\cap J_{Y_{0}}))\geq H_{d}(S/I)+q$. As shown in Equation A.6, the opposite inequality is always satisfied, so Equation A.3 follows. ∎
In the remainder of the section we prove different numerical criteria that are used in Section 6. 6.5 therein is in effect.
Proposition A.2.
Let $d\in{\mathbb{N}}$ satisfy $d\geq 6$. Then $s^{\prime\prime\prime}_{d}<2\cdot s_{d-3}$.
Proof.
The table below shows the inequality for $6\leq d\leq 8$.
$d=6$ $d=7$ $d=8$
$s^{\prime\prime\prime}_{d}$ $4$ $4$ $5$
$2\cdot s_{d-3}$ $6$ $8$ $10$
Next we show $s^{\prime\prime\prime}_{d+3}<2\cdot s_{d}$ for all $d\geq 6$. Let $d=6k+i$ where $1\leq k$ and $0\leq i\leq 5$. By Remark 6.6 we get
$s^{\prime\prime\prime}_{d+3}=3k+\left\lfloor\frac{i+3}{2}\right\rfloor+1\leq 3k+5\text{ since }i\leq 5.$
Furthermore, by Remark 6.6 (1) we have $s_{6k}=3k^{2}+3k+1\leq s_{d}$. Since $1\leq k$ the strict inequality holds and establishes
$s^{\prime\prime\prime}_{d}\leq 3k+5<2\cdot s_{6k}=2\cdot(3k^{2}+3k+1)\leq 2\cdot s_{d}\qed$
Lemma A.3.
Let $d\in N$ be such that $d\geq 6$. We state and prove all cases of the inequality Equation 6.7 used in the proof of Lemma 6.9.
  1. (1)
    $\frac{s^{\prime}_{d}}{2}\leq s^{\prime}_{d-1}$
  2. (2)
    $\frac{s^{\prime\prime}_{d}}{2}\leq s^{\prime\prime}_{d-2}$
  3. (3)
    $\frac{s^{\prime\prime\prime}_{d}}{2}\leq s^{\prime\prime\prime}_{d-3}$
Proof.
$(1)$ Let $d=6p+i$ where $p\geq 1$. Using Remark 6.6 we have
$s^{\prime}_{d}=\begin{cases}p&i=1\\ p+1&\text{otherwise}\end{cases}\qquad\text{ and }\qquad s^{\prime}_{d-1}=\begin{cases}p&i=0,2\\ p+1&\text{otherwise.}\end{cases}$
Hence it follows
$\frac{s^{\prime}_{d}}{2}\leq\frac{p+1}{2}\leq p\leq s^{\prime}_{d-1}.$
$(2)$ Let $d=3m+j$ where $m\geq 2$. Using Remark 6.6 we have
$s^{\prime\prime}_{d}=\left\lfloor\frac{d}{3}\right\rfloor+1\qquad\text{ and }\qquad s^{\prime\prime}_{d-2}=\begin{cases}m&j=0,1\\ m+1&j=2.\end{cases}$
Then
$\frac{s^{\prime\prime}_{d}}{2}=\frac{m+1}{2}\leq m\leq s^{\prime\prime}_{d-2}.$
$(3)$ Let $d=2s+t$ where $s\geq 3$. Recall from Remark 6.6 that
$s^{\prime\prime\prime}_{d}=\left\lfloor\frac{d}{2}\right\rfloor+1\qquad\text{ and }\qquad s^{\prime\prime}_{d-3}=\begin{cases}s-1&t=0\\ s&t=1.\end{cases}$
It follows that
$\frac{s^{\prime\prime\prime}_{d}}{2}=\cfrac{s+1}{2}\leq s-1\leq s^{\prime\prime\prime}_{d-3}.\qed$
Appendix B Secant varieties to the Veronese variety
Definition B.1.
Let $S$ be a polynomial ring. Define the $d$-th Veronese variety as $V^{d}=\operatorname{Proj}(k[S_{d}])$.
Remark B.2.
The Veronese variety can be embedded in (straight) projective space as follows: Let $S=k[x_{0},\ldots,x_{n}]$ and $N=s_{d}-1$. Let $m_{0},\ldots,m_{N}$ enumerate the monomials in $S_{d}$. Consider the $k$-algebra homomorphism
$\begin{split}\theta:k[y_{0},\ldots,y_{N}]&\to k[x_{0},\ldots,x_{n}]\\ y_{i}&\to m_{i}.\end{split}$ (B.1)
Set $R=k[y_{0},\ldots,y_{N}]$ to be the domain of $\theta$ and $I$ to be the kernel of $\theta$. Then Equation B.1 gives rise to an isomorphism $R/I\cong k[S_{d}]$ which in turn induces an isomorphism of projective varieties $V^{d}\cong\operatorname{Proj}(R/I)\subset\operatorname{Proj}(R)={\mathbb{P}}^{N}_{k}$.
Geometrically $\theta$ induces a map of projective varieties
$\psi:{\mathbb{P}}\dashrightarrow V^{d},\psi(p)=[m_{0}(p):\cdots:m_{N}(p)]$ (B.2)
defined on $U=\{p\in\operatorname{Proj}(S)\;|\;\theta(R_{+})\not\subseteq p\}$.
The main goal of this section is to establish the following proposition.
Theorem 7.5.
Let ${\mathbb{P}}={\mathbb{P}}(1,a_{1},\ldots,a_{n})$. For a fixed $d\geq a_{n}$ the following are equivalent:
  1. (1)
    if $X$ is a set of $r$ general double points, ${H_{d}(S/I_{X})=\min\{s_{d},r(n+1)\}}$
  2. (2)
    $\dim_{k}(\sigma_{r}(V^{d}))=\min\{s_{d}-1,r(n+1)-1\}$.
To show this we make use of tangent spaces which we now define.
Definition B.3.
Let $p$ be a point of an affine variety $U$ with defining ideal ${\mathfrak{m}}_{p}$. The Zariski tangent space to $U$ at $p$ is the vector space
$T_{p}(U)=\operatorname{Hom}_{k}({\mathfrak{m}}_{p}/{\mathfrak{m}}_{p}^{2},k).$
If $p$ is a point of a projective variety $X$, the projective tangent space to $X$ at $p$, denoted ${\mathbb{T}}_{p}(X)$ is the projectivization of the $T_{\widetilde{p}}\widetilde{X}$, where $\widetilde{X}$ is affine cone over $X$ and $\widetilde{p}$ is any point of the affine cone that corresponds to $p\in X$.
Lemma B.4.
Consider a map $g:{\mathbb{A}}_{k}^{n}\to{\mathbb{A}}_{k}^{N}$ given by $g(p)=(g_{1}(p),\ldots,g_{N}(p))$, where $g_{i}\in k[x_{1},\ldots,x_{n}]$. For any $p\in{\mathbb{A}}^{n}$ the induced map
${\rm d}g:T_{p}{\mathbb{A}}^{n}\to T_{g(p)}{\mathbb{A}}^{N}\text{ is represented by the Jacobian matrix }\begin{bmatrix}\frac{\partial g_{i}}{\partial x_{j}}\big{|}_{p}\end{bmatrix}_{\begin{subarray}{c}1\leq i\leq N\\ 1\leq j\leq n\end{subarray}}.$
We now analyze the projective tangent space to the Veronese variety $V^{d}$.
Proposition B.5.
Let ${\mathbb{P}}={\mathbb{P}}(1,a_{1},\ldots,a_{n})$, let $d\geq a_{n}$ be an integer and denote $N=s_{d}-1$. Using the notation introduced in Equation B.1 and Equation B.2, consider a point $p\in U_{0}={\mathbb{P}}\setminus V(x_{0})$ and set $U^{\prime}_{0}=V^{d}\cap({\mathbb{P}}_{k}^{N}\setminus V(y_{0}))$. Then
  1. (1)
    The inclusion $k[S_{d}]\hookrightarrow S$ induces an isomorphism between $U_{0}$ and $U_{0}^{\prime}$, both of which are isomorphic to ${\mathbb{A}}_{k}^{n}$. In particular, the points $p$ and $\psi(p)$ are smooth on ${\mathbb{P}}$ and $V^{d}$, respectively.
  2. (2)
    The point $\psi(p)$ is defined and the projective tangent space ${\mathbb{T}}_{\psi(p)}(V^{d})$ is given by
[…]
Proof.
(1) Let $S=k[x_{0},\ldots,x_{n}]$ with $\deg(x_{i})=a_{i}$. Recall that ${\mathbb{P}}=\operatorname{Spec}(S)$. Moreover the coordinate ring of $U_{0}$ is $S[x_{0}^{-1}]_{0}\cong k[x_{1},\ldots,x_{n}]$ since $\deg(x_{0})=1$. See the description of the coordinate rings of $U_{i}$ in Equation 2.3 for a justification of this isomorphism. Since the coordinate ring of $U_{0}$ is regular, all points of $U_{0}$ are smooth.
Now consider $V^{d}=\operatorname{Spec}(k[S_{d}])$. The following are monomials in $S_{d}$
$m_{0}=x_{0}^{d},m_{1}=x_{0}^{d-a_{1}}x_{1},\ldots,m_{n}=x_{0}^{d-a_{n}}x_{n},$
and every other monomial $m$ in $S_{d}$ can be written in the form
$m=\prod_{i=0}^{n}x_{i}^{e_{i}}=\frac{x_{0}^{e_{0}}\prod_{i=1}^{n}m_{i}^{e_{i}}}{x_{0}^{\sum_{i=1}^{n}e_{i}(d-a_{i})}}=\frac{x_{0}^{\sum_{i=0}^{n}e_{i}a_{i}}\prod_{i=1}^{n}m_{i}^{e_{i}}}{x_{0}^{\sum_{i=1}^{n}e_{i}d}}=\frac{m_{0}\prod_{i=1}^{n}m_{i}^{e_{i}}}{m_{0}^{\sum_{i=1}^{n}e_{i}}}.$
It follows that $k[S_{d},x_{0}^{-d}]_{0}\cong k[\frac{m_{1}}{m_{0}},\ldots,\frac{m_{n}}{m_{0}}]$. Since $\frac{m_{1}}{m_{0}},\ldots,\frac{m_{n}}{m_{0}}$ each contain a distinct variable $x_{i}$, they are algebraically independent. Thus the coordinate ring of $U^{\prime}_{0}$ is regular and we deduce that $V^{d}$ is smooth at $q$ whenever $q\in U^{\prime}_{0}$. In particular, the points $\psi(p)$ with $p\in U_{0}$ are smooth points of $V^{d}$.
The inclusion $\iota:k[S_{d}]\hookrightarrow S$ yields upon inverting powers of $x_{0}$ the identity map
$k\left[\frac{m_{1}}{m_{0}},\ldots,\frac{m_{n}}{m_{0}}\right]=k[S_{d},x_{0}^{-d}]\to S[x_{0}^{-1}]=k\left[\frac{x_{1}}{x_{0}},\ldots,\frac{x_{n}}{x_{0}}\right]$
since $m_{i}/m_{0}\mapsto x_{i}/x_{0}=m_{i}/m_{0}$ for $1\leq i\leq n$.
(2) The point $\psi(p)$ is defined since $m_{0}(p)\neq 0$.
Consider the affine cone $\widetilde{V^{d}}$ of $V^{d}$ and the regular map
$g:{\mathbb{A}}_{k}^{n+1}\to{\mathbb{A}}_{k}^{N+1},g(p)=(m_{0}(p),\ldots,m_{N}(p)).$
The map induced by $g$ on coordinate rings is denoted $\theta$ in Equation B.1.
Let ${\mathfrak{m}}$ denote the ideal of $p$ in ${\mathbb{A}}_{k}^{n+1}$, ${\mathfrak{m}}^{\prime}$ denote the ideal of $g(p)$ in ${\mathbb{A}}_{k}^{N+1}$, and ${\mathfrak{m}}^{\prime\prime}$ denote the ideal of $g(p)$ in $A^{\prime}=k[\widetilde{V^{d}}]$. Let $\pi:k[{\mathbb{A}}_{k}^{N+1}]\to A^{\prime}$ denote the canonical projection and $\iota:k[S_{d}]\to S=k[A^{n+1}]$ denote the inclusion. These maps satisfy the identity $\theta=\iota\circ\pi$. There are dual commutative diagrams as follows
[…]
By part (1), the horizontal map labeled $\iota$ and hence also its dual ${\rm d}\iota$ are isomorphisms. Since $\pi$ is surjective, ${\rm d}\pi$ is injective. Hence $T_{g(p)}\widetilde{V^{d}}$ is isomorphic to the image of ${\rm d}\pi$ which coincides with the image of ${\rm d}g$. By Lemma B.4, the differential ${\rm d}g:T_{p}{\mathbb{A}}_{k}^{n+1}\to T_{g(p)}{\mathbb{A}}_{k}^{N+1}$ is represented by the Jacobian matrix $J=\begin{bmatrix}\frac{\partial m_{i}}{\partial x_{j}}\end{bmatrix}_{\begin{subarray}{c}0\leq i\leq N\\ 0\leq j\leq n\end{subarray}}$. Therefore, $T_{\tilde{\psi(p)}}(\widetilde{V^{d}})$ can be identified with the column space of $J$ and its projectivization yields
${\mathbb{T}}_{\psi(p)}(V^{d})\cong{\mathbb{P}}\left\langle\left[\frac{\partial m_{0}}{\partial{x_{j}}}\big{|}_{p},\ldots,\frac{\partial m_{N}}{\partial{x_{j}}}\big{|}_{p}\right]\mid 0\leq j\leq n\right\rangle.$
In the following, we to denote by $\widetilde{V^{d}}$ the affine cone over $V^{d}$ and by $\widetilde{q}$ any affine point that corresponds to the projective point $q\in{\mathbb{P}}_{k}^{N}$. Proposition B.6 in standard graded case was first proven by Lasker in [Las04]. Our proof is an adaption from [BO08].
Proposition B.6 (Lasker’s proposition).
Let ${\mathbb{P}}={\mathbb{P}}(1,a_{1},\ldots,a_{n})$, let $d\geq a_{n}$ be an integer and set $N=s_{d}-1$. Let $p\in U_{0}={\mathbb{P}}\setminus V(x_{0})$. Then the orthogonal complement to the tangent space to $\widetilde{V^{d}}$ at $\widetilde{\psi(p)}$ consists of all hypersurfaces of degree $d$ singular at $p$. In detail, there is a vector space isomorphism
$[I_{p}^{(2)}]_{d}\cong T_{\widetilde{\psi(p)}}(\widetilde{V^{d}})^{\perp}.$
Proof.
By the Zariski-Nagata Theorem, $[I_{p}^{(2)}]_{d}$ is the vector space of homogeneous polynomials of degree $d$ whose first derivative vanishes on $p$
$(I_{p}^{(2)})_{d}=\left\{f\in S_{d}\;\big{|}\;\frac{\partial f}{\partial x_{i}}(p)=0\right\}=\left\{\displaystyle\sum_{i=0}^{N}\lambda_{i}m_{i}\;\big{|}\;\sum_{i=0}^{N}\lambda_{i}\frac{\partial m_{i}}{\partial x_{i}}\big{|}_{p}=0\right\}.$
Then for any $\underline{\lambda}\in k^{N+1}$ we have
$\begin{split}&\sum_{i=0}^{N}\lambda_{i}\frac{\partial m_{i}}{\partial x_{i}}\big{|}_{p}=0\\ \Leftrightarrow&(\lambda_{0},\ldots,\lambda_{N})\cdot\left(\frac{\partial m_{1}}{\partial x_{i}}\big{|}_{p},\cdots,\frac{\partial m_{N}}{\partial x_{i}}\big{|}_{p}\right)=0\\ \Leftrightarrow&\underline{\lambda}\in T_{\widetilde{\psi_{0}(p)}}(\widetilde{V^{d}})^{\perp}.\qed\end{split}$
We now define the $k$-th secant variety of a projective variety. This is a central tool in establishing Theorem 7.5. For a more thorough study of the secant variety see [Har92, Lecture 8].
Definition B.7 (Secant Variety).
Let $X$ be a projective variety. For any nonnegative integer $r$, the $r$-secant variety of $X$, denoted by $\sigma_{r}(X)$, is defined to be
$\sigma_{r}(X)=\overline{\bigcup_{P_{1},\dots,P_{r}\in X}\langle P_{1},\dots,P_{r}\rangle}^{\text{ Zariski closure}}.$
Note that $\sigma_{r}(V^{n}_{d})$ is an irreducible variety for all $r$. ([Fan90, Remark 1.2])
Lemma B.8 (First Terracini Lemma).
Let $Y$ be an a affine variety set, let $p_{1},\ldots,p_{r}\in Y$ be general points and $z\in\left<p_{1},\ldots,p_{r}\right>$ be a general point. Then
$T_{z}(\sigma_{r}(Y))=\left<T_{p_{1}}(Y),\ldots,T_{p_{r}}(Y)\right>.$
Proof.
See [BO08, Lemma 2.2]. ∎
Using arguments identical to [BO08] we prove Theorem 7.5.
Proof of Theorem 7.5.
Let $X=\{2p_{1},\ldots,2p_{r}\}$ be a collection of general double points in ${\mathbb{P}}$. Applying Proposition B.6 to get the second equality, we obtain
$[I_{X}]_{d}=\left[\bigcap_{i=1}^{r}I_{p_{i}}^{(2)}\right]_{d}=\displaystyle\bigcap_{i=1}^{r}\left(T_{\widetilde{\psi(p_{i})}}\widetilde{V^{d}}\right)^{\bot}=\left<T_{\widetilde{\psi(p_{1})}}\widetilde{V^{d}},\ldots,T_{\widetilde{\psi(p_{r})}}\widetilde{V^{d}}\right>^{\bot}.$
Therefore, using Lemma B.8, we obtain
$\dim([S/I_{X}]_{d})=\dim_{k}\left\langle T_{\widetilde{\psi(p_{1})}}\widetilde{V^{d}},\ldots,T_{\widetilde{\psi(p_{r})}}\widetilde{V^{d}}\right\rangle=\dim_{k}T_{z}(\sigma_{k}(\widetilde{V^{d}})).$
Finally using that $\sigma_{r}(\widetilde{V^{d}})=\widetilde{\sigma_{r}(V^{d})}$ and the irreducibility of $\widetilde{V^{d}}$, which implies irreducibility of $\sigma_{r}(\widetilde{V^{d}})$ for all $r$ [Har92, p. 144, Prop 11.24], we get for $z$ a general point of $\sigma_{r}(\widetilde{V^{d}})$
$\dim T_{z}(\sigma_{k}(\widetilde{V^{d}}))=\dim\sigma_{k}(\widetilde{V^{d}})=\dim(\sigma_{k}(V^{d}))+1.$
To summarize, we have obtained the identity
$\dim([S/I_{X}]_{d})=\dim(\sigma_{k}(V^{d}))+1,$
whence the claimed equivalence follows. ∎